### v1

# Functional Stochastic Gradient MCMC for Bayesian Neural Networks

Abstract
Classical variational inference for Bayesian neural networks (BNNs) in parameter space usually suffers from unresolved prior issues such as knowledge encoding intractability and pathological behaviors in deep networks, which could lead to an improper posterior inference. Hence, functional variational inference has been proposed recently to resolve these issues via stochastic process priors. Beyond variational inference, stochastic gradient Markov Chain Monte Carlo (SGMCMC) is another scalable and effective inference method for BNNs to asymptotically generate samples from true posterior by simulating a continuous dynamic. However, the existing SGMCMC methods only work in parametric space, which has the same issues of parameter-space variational inference, and extending the parameter-space dynamics to function-space dynamics is not a trivial undertaking. In this paper, we introduce a new functional SGMCMC scheme via newly designed diffusion dynamics, which can incorporate more informative functional priors. Moreover, we prove that the stationary distribution of these functional dynamics is the target posterior distribution over functions. We demonstrate better performance in both accuracy and uncertainty quantification of our functional SGMCMC on several tasks compared with naive SGMCMC and functional variational inference methods.
Machine Learning, ICML
1 Introduction
Compared to regular deep neural networks that rely on a single pattern of model parameters only, Bayesian neural networks (BNNs) have demonstrated the ability to provide a principled predictive uncertainty estimation and enhanced generalization performance through a model integration under posterior distribution over network weights (Neal, 1995; Gal et al., 2016; Wilson & Izmailov, 2020). With the continued development and success of Bayesian inference methods in the deep learning community, BNNs have been applied in a wide variety of domains and tasks, including prediction in astrophysics (Cranmer et al., 2021), medical diagnosis and diabetic detection (Filos et al., 2019; Band et al., 2022), and other safety-critical environments (Rudner et al., 2022).
Prior choosing is apparently crucial for BNNs. The existing common practice is i.i.d. Gaussian priors for all network weights, which is also named parameter-space prior. However, this parameter-space prior has been shown to have some problematic issues. For example, as network depth increases, the function samples of Gaussian prior over weights tend to be horizontal (Duvenaud et al., 2014; Matthews et al., 2018), and the prior distribution of unit outputs becomes more heavy-tailed with ReLU activation (Vladimirova et al., 2019; Tran et al., 2022), which will lead to improper posterior inference (Papamarkou et al., 2022). Moreover, the effects of this parameter space prior to the posterior inference remain an open question due to the complex network architecture (Fortuin et al., 2021; Wild et al., 2022), so it is difficult to transform the valid prior knowledge into corresponding network weights’ prior. Given these unresolved prior problems for parameters, Bayesian inference directly in function space has gained increasing interest in recent years since one can incorporate more informative stochastic processes as the prior. For example, Gaussian processes (GPs) which can easily encode the prior knowledge about the potential functions via kernel functions choosing have been widely used as the prior in function-space variational inference for BNNs (Sun et al., 2019; Ma & Hernández-Lobato, 2021; Rudner et al., 2022; Pielok et al., 2022) which has been shown to yield better prediction and improved uncertainty estimation performance than variational inference in parameter space.
For the posterior inference given prior, Markov Chain Monte Carlo (MCMC) is another valid method for BNNs apart from the commonly used variational inference (Blundell et al., 2015; Sun et al., 2019). Instead of optimizing a variational posterior by minimizing the distance (divergence) with the true posterior (Graves, 2011; Blundell et al., 2015; Gal & Ghahramani, 2016), MCMC approaches are regarded as the gold standard for Bayesian inference since they generate discrete samples from a continuous-time Markov chain whose stationary distribution is exactly the true posterior (Neal et al., 2011; Welling & Teh, 2011; Li et al., 2016). So, they do not need to make strong assumptions about the distributional form of the posterior (Alexos et al., 2022; Cobb & Jalaian, 2021). Even better, the computationally expensive MCMC methods became feasible for modern BNNs with large data following the introduction of their scalable stochastic gradient variants (SGMCMC) (Welling & Teh, 2011; Patterson & Teh, 2013; Chen et al., 2014; Zhang et al., 2020; Nemeth & Fearnhead, 2021). However, all existing MCMC methods for BNNs are still in parameter space, and extending them to function space is not a minor process because we need to design novel dynamics in function space and ensure the new dynamics would have the target posterior as the stationary distribution.
In this work, we propose a functional SGMCMC scheme based on the fundamental Itô lemma for stochastic calculus (Itô, 1951; Brzeźniak et al., 2008; Øksendal, 2013). The dynamics of BNNs in function space are designed by proposing a new potential energy function where we integrate the more informative functional prior, such as GPs, into the diffusion process. Furthermore, we prove that the stationary distribution of the new functional dynamics is the desired posterior distribution over functions. Our main contributions are as follows:
  • We propose a new functional SGMCMC approach for BNNs that can utilize the more informative functional prior to the posterior inference.
  • We prove that under our designed functional SGMCMC dynamics, the stationary distribution would be the target posterior distribution over functions of BNNs.
  • Our method shows improved predictive performance and uncertainty estimation in several benchmark tasks compared to parameter-space SGMCMC and parameter- and function-space variational inference methods.
2 Preliminaries
2.1 Bayesian Neural Networks
[…]
2.2 Stochastic Gradient MCMC for BNNs
Traditional MCMC methods, such as the random-walk Metropolis-Hastings (MH) algorithm (Metropolis et al., 1953), are easy to apply but suffer from a slow mixing rate and high correlation between samples. In contrast, the dynamics-based sampling methods utilizing gradients of log posterior density can provide a more efficient exploration of the posterior (Girolami & Calderhead, 2011). Also, their stochastic gradient variants (Welling & Teh, 2011; Chen et al., 2014; Baker et al., 2019), such as the stochastic gradient Langevin dynamics (SGLD) and the stochastic gradient Hamiltonian Monte Carlo (SGHMC), can scale to large datasets for practical applications.
SGLD. The objective of MCMC is to obtain random samples of the posterior distribution of model parameters (with the unnormalized form) $p(\mathbf{w}|\mathcal{D})\propto\exp(-U(\mathbf{w}))$, where $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})$ is the potential energy function. The idea of dynamics-based sampling for $p(\mathbf{w}|\mathcal{D})$ is to simulate a continuous-time stochastic process whose stationary distribution is the target posterior. Combining Langevin dynamics (Neal et al., 2011) with stochastic Robbins-Monro optimization algorithm (Robbins & Monro, 1951), Welling & Teh (2011) proposed SGLD for scalable MCMC algorithms. The approximated discretization transition rule for posterior samples is
$\mathbf{w}_{t+1}=\mathbf{w}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})+\sqrt{2\epsilon_{t}}\eta_{t},$ (1)
where $\epsilon_{t}$ is the step-size, $\eta_{t}$ is a standard Gaussian, and $\nabla\tilde{U}(\mathbf{w})$ uses mini-batch stochastic gradient for the likelihood term as $\nabla\tilde{U}(\mathbf{w})=-\frac{N}{n}\sum_{i=1}^{n}\nabla\log p(y_{i}|x_{i},\mathbf{w})-\nabla\log p(\mathbf{w})$ and $n\ll N$. With the decreasing step-size $\epsilon_{t}$, the stochastic gradient noise tends to zero faster than the injected Gaussian noise $\eta_{t}$, and the discretization error is eliminated, then the MH step can be ignored.
SGHMC. To enable a more efficient exploration in the state space, Hamiltonian dynamics (Duane et al., 1987; Neal, 2012) introduces a momentum variable $\mathbf{z}\in\mathbb{R}^{d}$, and the new joint distribution is $p(\mathbf{w}|\mathcal{D})p(\mathbf{z})\propto\exp(-H(\mathbf{w},\mathbf{z}))$ with corresponding energy function $H(\mathbf{w},\mathbf{z})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})-\log p(\mathbf{z})$, where $p(\mathbf{z})$ is commonly assumed as a zero-mean Gaussian $\mathcal{N}(\mathbf{z}|0,M)$. The sample rule in the full-batch Hamiltonian Monte Carlo is defined as following with an MH step (Neal, 2012; Cobb & Jalaian, 2021):
$\mathbf{w}_{t+1}$ $=\mathbf{w}_{t}+\epsilon_{t}M^{-1}\mathbf{z}_{t},$ (2)
$\mathbf{z}_{t+1}$ $=\mathbf{z}_{t}-\epsilon_{t}\nabla U\left(\mathbf{w}_{t}\right).$
Like SGLD, SGHMC uses stochastic gradient $\nabla\tilde{U}(\mathbf{w})$ for the sampling of momentum $\mathbf{z}$, also with an additional friction term $\epsilon_{t}M^{-1}\mathbf{z}_{t}$ and a Gaussian injected noise to guarantee the (marginal) stationary distribution of the Hamiltonian dynamics still be the target posterior distribution (Chen et al., 2014). The specific update rule for samples of $\mathbf{z}$ is as: $\mathbf{z}_{t+1}=\mathbf{z}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})-\epsilon_{t}CM^{-1}\mathbf{z}_{t}+\sqrt{2C\epsilon_{t}}\eta_{t}$, where $C>0$ is a scalar coefficient.
3 Functional Stochastic Gradient MCMC
Generally, dynamics-based MCMC approaches are derived from the Itô diffusion (Øksendal, 2003), which is similar to the Brownian motion of the particle subjected to a potential in fluid physics. In this section, to solve the parameter-space prior issues in existing SGMCMC methods, we propose a new functional SGMCMC scheme via a newly designed diffusion dynamics, which is capable of incorporating more informative functional priors. In Section 3.1, we first review the Itô diffusion for dynamics-based MCMC. Then we design a new functional Langevin dynamics directly for the function mapping defined by the BNNs. We prove that the stationary distribution of such functional Langevin dynamics is the target posterior distribution over functions. Furthermore, in Section 3.2, we design a functional Hamiltonian dynamics and the corresponding stochastic gradient sampling scheme for BNNs. We also prove it can guarantee the desired invariant distribution.
3.1 Functional SGLD for BNNs
An Itô diffusion in n-dimensional Euclidean space is defined as a stochastic differential equation (SDE) to describe the evolution of a continuous-time stochastic process $X:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}$ on a probability space $(\Omega,\Sigma,P)$. The general form of an Itô diffusion driven by the standard Wiener process (Brownian motion) is as follows:
$\mathrm{d}X_{t}=\mu(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},$ (3)
where $X_{t}$ represents the state of the stochastic process at time $t$, $B\in\mathbb{R}^{n}$ is an n-dimensional Brownian motion, $\mu(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a vector field denoting the deterministic drift term and matrix field $\sigma(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ is known as the diffusion coefficient for $X_{t}$. Both $\mu(\cdot)$ and $\sigma(\cdot)$ are assumed to satisfy the usual Lipschitz continuity condition (Ghosh, 2010). And $\mathrm{d}B_{t}$ is the increment of a Wiener process distributed as $\mathrm{d}B_{t}\sim\mathcal{N}(0,\mathrm{d}t\cdot\mathbf{I}_{n})$. Let $p(x,t)$ be the probability density of random variable $X_{t}$, and the corresponding evolution of this probability density of the diffusion process can be described by the Fokker–Planck equation (Risken, 1996), which is the key to derive the dynamics-based MCMC (see Appendix B).
Naturally, the idea of existing dynamics-based SGMCMC methods for BNNs is to simulate a continuous diffusion of network weight whose stationary distribution is the target posterior $p(\mathbf{w}|\mathcal{D})$ with appropriately designed drift term and diffusion coefficient. To incorporate informative functional prior into the posterior inference process, we now directly consider the dynamics of function mapping $f(\cdot;\mathbf{w})$ defined by BNNs. According to Itô Lemma (Itô, 1951), $f(\cdot;\mathbf{w})$ is itself an Itô diffusion (Brzeźniak et al., 2008) (see Appendix A). That is, for exiting SGMCMC methods, the stochastic process $X$ simulated in Itô diffusion is usually for network weight or its augmentation, which we now lift to the stochastic function mapping of BNNs directly as
$\mathrm{d}f_{t}=\mu(f_{t})\mathrm{d}t+\sigma(f_{t})\mathrm{d}B_{t}.$ (4)
In a probability space $(\Omega,\mathcal{F},P)$ (separable and complete Polish space), suppose $p(f)$ is the prior distribution over functions, such as the classical Gaussian process (GP) prior that can easily encode prior knowledge about function properties. $p(\mathbf{Y_{\mathcal{D}}}|f(\mathbf{X_{\mathcal{D}}}))$ is the likelihood function with observation $\mathcal{D}=\left\{\mathbf{X_{\mathcal{D}}},\mathbf{Y_{\mathcal{D}}}\right\}$ and $p(f|\mathcal{D})$ is the posterior over functions. For brevity, we abbreviate them as $p(f)$, $p(\mathcal{D}|f)$ and $p(f|\mathcal{D})$, respectively. Let $U(f)=-\log p(\mathcal{D}|f)-\log\frac{\mathrm{d}p}{\mathrm{d}r_{\mathcal{H,F}}}(f))$ be the potential energy functional, where $r_{\mathcal{H,F}}$ is the corresponding Hausdorff / Riemannian measure, and $\mu(f)=-\nabla_{f}U(f),\sigma(f)=\sqrt{2}\cdot\mathbf{I}_{\infty}$ for the function mapping $f$ of BNNs, we have the following functional Langevin dynamics:
$\mathrm{d}f_{t}=-\nabla_{f}U(f_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$ (5)
and it is easy to verify that the stationary distribution for $p(f,t)$ is exactly the target posterior distribution over functions as $\pi(f)=\exp(-U(f))=p(f|\mathcal{D})$.
However, even for the finite input case, it is intractable to discretize and sample the above diffusion of function mapping in practice directly. Considering that a function mapping $f(\cdot;\mathbf{w})$ of BNNs is completely determined by the network weights $\mathbf{w}$ (assuming the network architecture is fixed), we solve this issue by transforming the functional Langevin dynamics to a weight-space dynamics of which stationary distribution is still $p(f|\mathcal{D})$. We firstly design the following specific Riemannian functional Langevin dynamics for $f(\cdot;\mathbf{w})$ as follows:
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=$ $\mu(f_{t}(\cdot;\mathbf{w}))\mathrm{d}t+\sigma(f_{t}(\cdot;\mathbf{w}))\mathrm{d}B_{t}$ (6)
$=$ $\left[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f\right]\mathrm{d}t$
$+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t},$
where $\nabla_{\mathbf{w}}f$ is the gradient of $f(\cdot;\mathbf{w})$ w.r.t. weight $\mathbf{w}$, $H_{\mathbf{w}}f$ is the Hessian matrix of $f(\cdot;\mathbf{w})$ w.r.t. $\mathbf{w}$. The stationary distribution of this dynamics is guaranteed to be $p(f|\mathcal{D})$ by the following Proposition.
Proposition 3.1.
The stationary distribution of the functional Langevin dynamics defined in Equation 6 is the target posterior distribution over functions $p(f|\mathcal{D})$.
Proof.
The proof is given in Appendix B. ∎
Secondly, we achieve the function-space to weight-space transformation based on the relationship between the dynamics of a random variable with the dynamics of its function. According to the Itô Lemma (Itô, 1951), given the general diffusion framework for weights as $\mathrm{d}\mathbf{w}_{t}=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$, the differential of the function mapping $f(\cdot;\mathbf{w})$ can be denoted as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ (7)
$=(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}\mathbf{w}_{t}+\frac{1}{2}(\mathrm{d}\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\mathrm{d}\mathbf{w}_{t}$
$=\left((\nabla_{\mathbf{w}}f)^{T}\mu(\mathbf{w}_{t})+\frac{1}{2}Tr[\sigma(\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\sigma(\mathbf{w}_{t})]\right)\mathrm{d}t$
$+(\nabla_{\mathbf{w}}f)^{T}\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}.$
Given our defined functional dynamics in Equation 6 and the above relationship, the corresponding diffusion process of network weights in parameter space under our functional Langevin dynamics can be derived as
$\mathrm{d}\mathbf{w}_{t}$ $=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$ (8)
$=\nabla_{\mathbf{w}}f[\nabla_{f}(\log p(\mathcal{D}|f_{t})+\log\frac{\mathrm{d}p_{0}}{\mathrm{d}r_{\mathcal{H,F}}}(f))]\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$
where the corresponding drift term $\mu(\mathbf{w}_{t})=-\nabla_{\mathbf{w}}U(f_{t})$, and the diffusion coefficient $\sigma(\mathbf{w}_{t})=\sqrt{2}\cdot\mathbf{I}_{d}$. Then, the function samples of the target posterior $p(f|\mathcal{D})$ as the stationary distribution of the functional Langevin dynamics can be obtained by discretizing and sampling the diffusion process of weights in Equation 8.
In practice, the stochastic gradient variant drift term of SDE in Equation 8 is denoted as $-\nabla_{\mathbf{w}}\tilde{U}(f_{t})$, which uses the minibatch version for the likelihood term. For the gradient of the log prior distribution over functions, we firstly assume an approximated GP prior for BNNs, since in the limit of infinite width, the prior measure of BNNs converges to a GP with NNGP kernel (Neal, 1995; Lee et al., 2017; Matthews et al., 2018; Lee et al., 2019). Due to the infinite-dimensional property of functional distribution, we solve the gradient of functional prior on finite measurement points $\mathbf{X}_{\mathcal{M}}\stackrel{{\scriptstyle\text{ dif }}}{{=}}\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\right]^{\mathrm{T}}$ as $\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f^{\mathbf{X}_{\mathcal{M}}})$, where $f^{\mathbf{X}_{\mathcal{M}}}$ are corresponding function values evaluated at $\mathbf{X}_{\mathcal{M}}$. Then the prior $p(f^{\mathbf{X}_{\mathcal{M}}})$ is a multivariate Gaussian that could be solved analytically for the gradient. The specific discretization update rule for weight samples under functional stochastic gradient Langevin dynamics (fSGLD) is
$\mathbf{w}_{t+1}=$ $\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\Big{[}-\frac{N}{n}\sum_{i=1}^{n}\nabla_{f}\log p(y_{i}|f_{t}(x_{i};\mathbf{w}_{t}))-$ (9)
$\hskip 62.59596pt\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f_{t}^{\mathbf{X}_{\mathcal{M}}})\Big{]}+\sqrt{2\epsilon_{t}}\eta_{t}.$
Similar to the naive SGLD, with the discretization step-size $\epsilon_{t}$ decays to zero, the stationary distribution for functional $p(f,t)$ of fSGLD continues to be the target posterior $p(f|\mathcal{D})$ over functions.
Note that although the function samples of posterior $p(f|\mathcal{D})$ are drawn by the discretization process for corresponding diffusion of weights in Equation 9, our fSGLD is different from the existing SGLD in parameter space, which means the stationary distributions from two dynamics would be different as well. Compared to the naive parameter-space SGLD in Equation 1 controlled only by the limited parameter information, the discretization update rule for weight samples in Equation 9 under our fSGLD scheme contains information about the function mapping, which is important for accurate posterior inference. The pseudocode for fSGLD is shown in Algorithm 1 in Appendix C.
3.2 Functional SGHMC for BNNs
For the naive Hamiltonian dynamics, $X=(\mathbf{w},\mathbf{z})$ is the augmentation of model parameters. We now let $X=(f,g)$ and define the Hamiltonian dynamics in function space as $H(f,g)=U(f)-\log p(g)$, where $p(g)$ is the functional auxiliary distribution induced by the distribution of auxiliary variable $\mathbf{z}$. We design the functional Hamiltonian dynamics for $f(\cdot;\mathbf{w})$ and $g(\cdot;\mathbf{z})$ as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})=-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (10)
$\mathrm{d}g_{t}(\cdot;\mathbf{z})=-(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f\nabla_{f}U(f_{t})\mathrm{d}t.$
Table 1: Drift term and diffusion coefficient for naive SGMCMC and functional SGMCMC algorithms.
Dynamics $\mu(X)$ $\sigma(X)$
SGLD $-\nabla\tilde{U}(\mathbf{w})$ $\sqrt{2}\cdot\mathbf{I}_{d}$
fSGLD $-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}\tilde{U}(f)+H_{\mathbf{w}}f$ $\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}$
SGHMC $-\left(\begin{array}[]{cc}\mathbf{0}&-\mathbf{I}_{d}\\ \mathbf{I}_{d}&C\end{array}\right)\left(\begin{array}[]{cc}\nabla\tilde{U}(\mathbf{w})\\ M^{-1}\mathbf{z}\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}\end{array}\right)$
fSGHMC $-\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)\left(\begin{array}[]{c}\nabla_{f}\tilde{U}(f)\\ -\nabla_{g}\log p(g)\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g)^{T}\end{array}\right)$
Proposition 3.2.
The (marginal) stationary distribution of the functional Hamiltonian dynamics defined in Equation 10 is the target posterior distribution over functions $p(f|\mathcal{D})$.
Proof.
The proof is given in Appendix B. ∎
Similar to the fSGLD, we also transform the above function-space Hamiltonian dynamics to its corresponding weight-space dynamics of $\mathbf{w}$ and $\mathbf{z}$ as follows:
$\mathrm{d}\mathbf{w}_{t}$ $=\frac{\partial H}{\partial\mathbf{z}}=-\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (11)
$\mathrm{d}\mathbf{z}_{t}$ $=-\frac{\partial H}{\partial\mathbf{w}}=-\nabla_{\mathbf{w}}f\cdot\nabla_{f}U(f_{t})\mathrm{d}t,$
where the random diffusion term is 0 for the full-batch version.
For the stochastic gradient version, the discretization update rule for samples of $\mathbf{w}$ and $\mathbf{z}$ under the functional stochastic gradient Hamiltonian dynamics (fSGHMC) is as follows:
$\mathbf{w}_{t+1}=$ $~{}\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})$ (12)
$\mathbf{z}_{t+1}=$ $~{}\mathbf{z}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\cdot\nabla_{f}\tilde{U}(f_{t})$
$\hskip 14.22636pt+\epsilon_{t}C\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})+\sqrt{2C\epsilon_{t}}\eta_{t}.$
The stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$ (see Appendix B). Like the discussion in the last section, the diffusion process for $\mathbf{w}$ and $\mathbf{z}$ under fSGHMC scheme in Equation 12 containing abundant functional information is obviously different from the naive parameter-space SGHMC in Equation 2. The pseudocode for fSGHMC is shown in Algorithm 2 in Appendix C.
For our functional SGMCMC, the predictive distribution is obtained from the following integration process:
$p(y^{*}|x^{*},\mathcal{D})$ $=\int p(y^{*}|f(x^{*}))p(f|\mathcal{D})\mathrm{d}f$ (13)
$\approx\frac{1}{S}\sum_{j=1}^{S}p(y^{*}|f(x^{*};\mathbf{w}^{(j)})),$
where function draws of posterior over functions are generated from corresponding samples of weight $\mathbf{w}^{(j)}(j=1,2,...,S)$ using fSGLD or fSGHMC.
The comparisons of the definitions for the drift term and the diffusion coefficient of popular exiting SGMCMC algorithms and our functional fSGLD and fSGHMC are summarized in Table 1.
4 Related Work
Functional Variational Inference for BNNs. With unresolved concerns about parameter-space weight prior to posterior approximation in BNNs, Bayesian inference in function space has attracted increasing attention. In order to specify meaningful functional prior like Gaussian processes, Sun et al. (2019) proposed a kind of functional variational Bayesian neural networks by maximizing the functional evidence lower bound (ELBO) to optimize the Kullback–Leibler (KL) divergence between the posterior processes and variational posterior over functions explicitly. They solved the KL divergence between infinite-dimensional stochastic processes on a finite marginal measurement set and used the Spectral Stein gradient estimator for implicit functional prior and variational posterior over functions of BNNs (Shi et al., 2018). However, Burt et al. (2020) pointed out some limitations of employing KL divergence in infinite-dimensional function space. The variational posterior should be absolutely continuous with respect to the prior measure to guarantee the existence of the Radon-Nikodym derivatives involved in the definition of functional KL divergence (Matthews et al., 2016; Wild et al., 2022). Otherwise, for example, the KL divergence between distributions over functions generated from two BNNs with different network architectures is ill-defined and could be infinite (Ma & Hernández-Lobato, 2021). To avoid this issue, Rudner et al. (2022) proposed to approximate the distributions over functions of a BNN prior and variational posterior that share the same network structures as two Gaussian distributions via the linearization of their mean parameters and derived a tractable and well-defined functional variational objective for BNNs. Considering the potential weaknesses of KL divergence for stochastic processes, Tran et al. (2022) proposed to match a BNN prior to a GP prior by minimizing the 1-Wasserstein distance (Kantorovich, 1960) to obtain a more interpretable functional prior for BNNs.
Markov Chain Monte Carlo for BNNs. Papamarkou et al. (2022) summarized the challenges in applying MCMC approaches to posterior inference for BNNs in four aspects: computational cost, model structure, weight symmetries, and prior specification. MCMC methods have not been broadly adopted in posterior inference for large modern BNNs until the introduction of scalable Monte Carlo algorithms, such as the stochastic gradient MCMC (Welling & Teh, 2011; Chen et al., 2014; Ahn et al., 2014; Garriga-Alonso & Fortuin, 2021). Ma et al. (2015) designed a unifying framework for stochastic gradient MCMC whereby all past dynamics-based samplers and their minibatch variants can be cast into it. For the multimodal properties of posterior, Zhang et al. (2020) proposed a cyclical step size schedule in stochastic gradient MCMC for BNNs to improve the sampling efficiency by which they proved the non-asymptotic convergence of their method. On the other hand, Cobb & Jalaian (2021) emphasized the limitations of the use of stochastic gradients in HMC (Shahbaba et al., 2014), and proposed a scaling full HMC for BNNs by a symmetric splitting integration scheme. And Izmailov et al. (2021) also used full-batch HMC to investigate the foundational properties of posterior for BNNs. Alexos et al. (2022) proposed a structured stochastic gradient MCMC that combines structure variational inference and MCMC to improve the slow mixing in high dimensional BNNs.
(a) fSGLD
(b) fSGHMC
(c) FBNN
(d) SGLD
(e) SGHMC
(f) IFBNN
Figure 1: 1-D extrapolation example. The green line is the ground true function, and the blue lines correspond to the mean of samples from posterior predictions. Black dots denote 20 training points; shadow areas represent the predictive standard deviations.
5 Experiments
We evaluate the predictive performance and uncertainty quantification of our functional stochastic gradient MCMC methods on several benchmark tasks, including a synthetic extrapolation example, multivariate regressions on UCI datasets, image classification tasks, and contextual bandits. We compare fSGLD and fSGHMC with naive SGLD, SGHMC, and competing functional variational inference methods to illustrate the sampling efficiency of our functional stochastic gradient MCMC methods.
5.1 Extrapolation on Synthetic Data
To evaluate the fitting ability and uncertainty estimation of our method, we first consider a 1-D oscillation curve extrapolation on a synthetic dataset. For input, we randomly sampled half of 20 observation points from Uniform$(-0.75,-0.25)$, and the other half are from Uniform$(0.25,0.75)$. The output is modeled through the polynomial function: $y=\sin(3\pi x)+0.3\cos(9\pi x)+0.5\sin(7\pi x)+\epsilon$ with noise $\epsilon\sim\mathcal{N}(0,0.5^{2})$. We compared our fSGLD and f SGHMC with naive SGLD, SGHMC, and two representative functional variational methods: FBNN (Sun et al., 2019) and IFBNN (Wu et al., 2023). The model is a fully-connected neural network with two hidden layers. For our functional fSGLD, fSGHMC, and the two functional variational inference methods, we use the same functional GP prior with the RBF kernel pre-trained on the input dataset. There are 40 measurement points for the approximate estimation of the gradient of log functional prior in fSGLD and fSGHMC, which are randomly sampled from training data and an additional 40 inducing points from Uniform $(-1,1)$. For all sampling methods, we use 2000 burn-in iterations and 8000 iterations for 80 samples (to reduce correlations between samples, we draw separated samples every 100 epochs). Functional variational inference methods are trained for 10000 epochs for fair comparison. Results are shown in Figure 1. Figure 1(a) and Figure 1(b) show that fSGLD and fSGHMC can recover the key characteristic of the curve in the observation range and provide reasonable uncertainty estimations in the unseen region. fSGHMC performs slightly better than fSGLD. In contrast, for naive SGMCMC methods, Figure 1(d) illustrates that SGLD underestimates the uncertainty in non-observation areas, especially evident in the leftmost non-observed $[-1,-0.75]$ interval. Meanwhile, Figure 1(e) shows that SGHMC suffers from some unreasonable manifestations of the uncertainty estimation, such as the undeserved uncertainty expansion in the interval $[-0.4,-0.25]$ of the input data range. On the other hand, the rightmost column shows results from two functional variational methods. We found that FBNN in Figure 1(c) is unable to fit the target curve in the training data range $[-0.75,0.25]$ and performs poorly for uncertainty quantification. Figure 1(f) shows that IFBNN also severely underestimates the uncertainty, which is consistent with the findings of Ormerod & Wand (2010) and Zhang et al. (2018) that variational approximation typically underestimates the posterior variance.
5.2 UCI Regression
In this experiment, we evaluate our method for multivariate regression tasks on 7 benchmark UCI datasets: Yacht, Boston, Concrete, Energy, Wine, Kin8nm and Protein. We use a 2- hidden-layer fully connected neural network. For all sampling methods, we run 500 iterations for the burn-in stage and collect 15 samples. Performance is measured via average root mean square error (RMSE) and the test negative log-likelihood (NLL). For each dataset, we construct 10 random 90-to-10 train-test splits and report the mean and standard deviation of RMSE and NLL over the splits in Table 2 and Table 3, respectively. fSGHMC outperforms all baselines for RMSE on all datasets, and fSGLD also achieves better accuracy values than naive SGMCMC methods and functional variational inference, demonstrating superior performance of our functional SGMCMC in predictive accuracy. Moreover, fSGHMC and fSGLD obtain better NLL results on most of the datasets (5/7), showing that our method is also very competitive in uncertainty quantification.
Table 2: The table shows the results of average RMSE for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
RMSE
SGLD fSGLD SGHMC fSGHMC FBNN IFBNN
Yacht 1.09 $\pm$ 0.10 0.41 $\pm$ 0.10 1.19 $\pm$ 0.10 0.25 $\pm$ 0.13 1.52 $\pm$ 0.08 1.24 $\pm$ 0.10
Boston 1.23 $\pm$ 0.06 0.36 $\pm$ 0.09 1.31 $\pm$ 0.06 0.24 $\pm$ 0.10 1.68 $\pm$ 0.12 1.44 $\pm$ 0.09
Concrete 1.10 $\pm$ 0.07 0.45 $\pm$ 0.04 1.15 $\pm$ 0.06 0.28 $\pm$ 0.05 1.27 $\pm$ 0.05 1.07 $\pm$ 0.07
Energy 1.04 $\pm$ 0.06 0.24 $\pm$ 0.03 1.07 $\pm$ 0.08 0.18 $\pm$ 0.03 1.35 $\pm$ 0.06 1.19 $\pm$ 0.05
Wine 1.08 $\pm$ 0.09 0.71 $\pm$ 0.05 1.14 $\pm$ 0.09 0.62 $\pm$ 0.04 1.53 $\pm$ 0.05 1.21 $\pm$ 0.05
Kin8nm 1.20 $\pm$ 0.02 0.74 $\pm$ 0.02 1.26 $\pm$ 0.03 0.37 $\pm$ 0.01 1.45 $\pm$ 0.07 1.12 $\pm$ 0.02
Protein 1.12 $\pm$ 0.01 0.84 $\pm$ 0.01 1.21 $\pm$ 0.01 0.83 $\pm$ 0.01 1.50 $\pm$ 0.03 1.16 $\pm$ 0.01
Table 3: The table shows the results of average NLL for multivariate regression on UCI datasets. We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
NLL
SGLD fSGLD SGHMC fSGHMC FBNN IFBNN
Yacht -0.37 $\pm$ 0.08 -2.46 $\pm$ 0.28 -2.06 $\pm$ 0.10 -3.37 $\pm$ 0.69 -0.77 $\pm$ 0.86 -1.25 $\pm$ 1.21
Boston -0.60 $\pm$ 0.11 -2.20 $\pm$ 0.20 -2.25 $\pm$ 0.08 -3.26 $\pm$0.33 -1.19 $\pm$ 0.76 0.32 $\pm$ 0.30
Concrete -0.54 $\pm$ 0.05 -1.63 $\pm$ 0.14 -2.51 $\pm$ 0.08 -2.74 $\pm$0.18 -1.00 $\pm$ 0.52 -0.39 $\pm$ 0.33
Energy -0.58 $\pm$ 0.09 -2.25 $\pm$ 0.25 -2.39 $\pm$ 0.09 -4.29 $\pm$0.23 -2.14 $\pm$ 0.47 -1.78 $\pm$ 0.36
Wine -0.90 $\pm$ 0.12 -2.53 $\pm$ 0.15 -2.50 $\pm$ 0.10 -2.34 $\pm$ 0.25 0.52 $\pm$ 0.14 0.26 $\pm$ 0.15
Kin8nm -0.60 $\pm$ 0.04 -2.24 $\pm$ 0.09 -2.29 $\pm$ 0.04 -2.00 $\pm$0.12 -2.45 $\pm$ 0.62 -1.01 $\pm$ 0.14
Protein -0.60 $\pm$ 0.01 -2.10 $\pm$ 0.08 -2.20 $\pm$ 0.01 -2.02 $\pm$0.06 -1.49 $\pm$ 0.24 -2.13 $\pm$ 0.34
Table 4: Image classification and OOD detection performance.
MNIST FMNIST CIFAR10
Model Test Error AUC Test Error AUC Test Error AUC
SGLD 2.73 $\pm$ 0.00 0.859 $\pm$ 0.06 13.92 $\pm$ 0.01 0.633 $\pm$ 0.13 49.56 $\pm$ 0.03 0.669 $\pm$ 0.04
fSGLD 1.89 $\pm$ 0.00 0.960 $\pm$ 0.02 11.66 $\pm$ 0.01 0.849 $\pm$ 0.03 33.57 $\pm$ 0.02 0.672 $\pm$ 0.03
SGHMC 2.97 $\pm$ 0.00 0.853 $\pm$ 0.05 13.72 $\pm$ 0.01 0.589 $\pm$ 0.07 48.13 $\pm$ 0.04 0.639 $\pm$ 0.06
fSGHMC 1.64 $\pm$ 0.00 0.931 $\pm$ 0.02 11.91 $\pm$ 0.00 0.880 $\pm$ 0.03 31.99 $\pm$ 0.01 0.717 $\pm$ 0.03
FBNN 3.99 $\pm$ 0.00 0.801 $\pm$ 0.03 14.36 $\pm$ 0.00 0.814 $\pm$ 0.02 53.71 $\pm$ 0.01 0.612 $\pm$ 0.03
IFBNN 3.64 $\pm$ 0.00 0.949 $\pm$ 0.03 14.15 $\pm$ 0.00 0.838 $\pm$ 0.01 53.38 $\pm$ 0.01 0.616 $\pm$ 0.03
(a) p = 0.4
(b) p = 0.5
(c) p = 0.6
Figure 2: Comparisons of cumulative regrets of fSGLD, SGLD, FBNN, IFBNN, BBB for contextual bandit task on the Mushroom dataset. Lower represents better performance.
5.3 Imagine Classification and OOD Detection
In this section, we investigate the scalability of functional SGMCMC on image classification tasks with high-dimensional inputs. We evaluate the in-distribution predictive performance and out-of-distribution (OOD) detection ability on MNIST (LeCun et al., 2010), FashionMNIST(Xiao et al., 2017) and CIFAR-10(Krizhevsky et al., 2009). We compare our fSGLD and fSGHMC with naive SGLD, SGHMC, functional variational FBNN and IFBNN. We use ResNet-18 architecture (He et al., 2016) for all methods on these three datasets, and the batch size is 125. For all sampling methods, we run for 60 burn-in iterations and collect 10 samples on the MNIST dataset, 100 burn-in iterations and collect 10 samples on FashionMNIST, 80 burn-in iterations and draw 10 samples on CIFAR10. The functional GP prior with the RBF kernel is pre-trained on the test dataset for all functional methods. We randomly sample 40 points from training data in every iteration as the measurement points for the gradient estimation of log functional prior distribution in fSGLD and fSGHMC. We report the test classification errors for predictive performance and the area under the curve (AUC) of OOD detection pairs FashionMNIST/MNIST, MNIST/FashionMNIST and CIFAR10/SVNH based on predictive entropies in Table 4. Our fSGLD and fSGHMC outperform all other methods for both predictive accuracy and OOD detection.
5.4 Contextual Bandits
In many downstream tasks, such as online sequential decision-making scenarios in contextual bandits problems, reliable uncertainty estimation is important to guide the exploration-exploitation process. In such problems, an agent interacts with an unknown environment repeatedly and chooses to take an optimal action in each round of interaction. Thompson sampling (Thompson, 1933; Russo & Van Roy, 2016) is a widely used algorithm for exploration strategies in contextual bandits, it (i) first samples a current posterior to obtain a model configuration; (ii) then adaptively chooses an optimal action based on the current context under the sampled model configuration and observes the corresponding reward; (iii) updates the posterior based on the context, action and reward tuples in this interaction round.
This section evaluates the ability to guide exploration on the UCI Mushroom dataset with 8124 instances. In each instance, the mushroom has 22 features and is labeled as edible or poisonous. An agent can observe mushroom features as the context in each interaction and choose to eat or reject a mushroom. We compare our fSGLD with several baselines, including naive SGLD, functional variational FBNN and IFBNN, parametric-space variational method Bayes By Backprop (BBB) (Blundell et al., 2015). We follow the basic settings by Blundell et al. (2015) and consider three different reward patterns: if the agent eats an edible mushroom, it receives a reward of 5 and a reward of 0 if the agent chooses to reject. On the other hand, if the agent eats a poisonous mushroom, it receives a reward of -35 with probabilities of 0.4, 0.5, and 0.6, respectively, for three different patterns. If the agent chooses to reject a poisonous mushroom, it receives a reward of 0. Suppose an oracle always chooses to eat an edible mushroom and reject the poisonous mushroom. The cumulative regrets with respect to the reward achieved by the oracle measure the exploration-exploitation ability of an agent. We run 10000 iterations for all methods and the minibatch size is 64. Figure 2 shows the cumulative regrets for all methods in three reward patterns. Our fSGLD consistently outperforms other inference methods, indicating that our functional SGMCMC can provide reliable uncertainty estimation in decision-making tasks.
6 Discussion and Conclusions
In this paper, we have lifted the Itô diffusion of network weights in BNNs to function space and proposed a new functional SGMCMC scheme for BNNs. We have designed the functional Langevin dynamics and Hamiltonian dynamics for the function mapping defined by BNNs directly, which can incorporate informative functional prior into posterior inference. Moreover, we proved that the stationary distribution of our functional SGMCMC is the posterior distribution over functions. Empirically, we demonstrated that our functional SGMCMC approaches can leverage functional priors to yield high predictive performance and principled uncertainty quantification on several benchmark tasks.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Further Background
Itô Lemma (Itô, 1951) is a fundamental result in stochastic calculus to find the differential of a function of a stochastic process, which serves as the stochastic calculus counterpart of the chain rule. For an Itô diffusion, let $f(X_{t})$ be an arbitrary twice differentiable scalar function of real variables $X_{t}$, then the differential of $f(X_{t})$ can be derived from the Taylor series expansion of the function as
$\mathrm{d}f(X_{t})=$ $\left(\mu(X_{t})\frac{\partial f}{\partial x}+\frac{\sigma^{2}(X_{t})}{2}\frac{\partial^{2}f}{\partial x^{2}}\right)\mathrm{d}t+\sigma(X_{t})\frac{\partial f}{\partial x}\mathrm{d}B_{t},$ (14)
which implies that $f(X_{t})$ is itself an Itô diffusion (Brzeźniak et al., 2008).
Appendix B Theoretical Proof
The Fokker–Planck equation (Risken, 1996) for the corresponding evolution of the probability density $p(x,t)$ of a diffusion process $X_{t}$ is as
$\frac{\partial p(x,t)}{\partial t}=-\sum_{i}\frac{\partial}{\partial x_{i}}\left[\mu_{i}(X_{t})p(x,t)\right]+\sum_{i,j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\left[D_{i,j}(x,t)p(x,t)\right],$ (15)
where $D_{i,j}(x,t)=\frac{1}{2}\sigma(X_{t})\sigma(X_{t})^{T}$. When the Fokker–Planck equation is equal to zero, it means that the diffusion process has reached a stationary state and $p(x,t)$ is equal to the stationary distribution $\pi(X)$. Then, simulating this diffusion process is equivalent to sampling from the stationary distribution.
The proof of Proposition 3.1 and Proposition 3.2 will be concise based on the general framework for dynamics-based sampling methods proposed by Ma et al. (2015). They proved that if $\mu(X_{t})$ and $\sigma(X_{t})$ are restricted to the following form
$\mu(X)=$ $-[\mathbf{D}(X)+\mathbf{Q}(X)]\nabla G(X)+\Gamma(X),$ (16)
$\Gamma(X)=$ $\sum\frac{\partial}{\partial X}(\mathbf{D}_{ij}(X)+\mathbf{Q}_{ij}(X)),$
$\sigma(X)=$ $\sqrt{2\mathbf{D}(X)},$
where $G(\cdot)$ is the potential energy function, $\mathbf{D}(\cdot)$ is a positive semidefinite matrix, and $\mathbf{Q}(\cdot)$ is a skew-symmetric curl matrix, then the Fokker–Planck equation can be transformed into a more compact form (Yin & Ao, 2006; Shi et al., 2012):
$\frac{\partial p(x,t)}{\partial t}=\nabla^{T}\cdot([\mathbf{D}(X)+\mathbf{Q}(X)][p(x,t)\nabla G(X)+\nabla p(x,t)]).$ (17)
Then, it is easy to verify that the stationary distribution of the diffusion process is exactly as $\pi(X)\propto\exp(-G(X))$.
For the proof of Proposition 3.1, it is clear that the functional Langevin dynamics in Equation 6 can be cast into the above general framework as
$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ $=[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f]\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t}$ (18)
$=\left(-[\mathbf{D}(f)+\mathbf{Q}(f)]\nabla_{f}U(f_{t})+\Gamma(f)\right)\mathrm{d}t+\sqrt{2\mathbf{D}(f)}\mathrm{d}B_{t},$
where $\nabla_{f}U(f)=-\nabla_{f}\log p(\mathcal{D}|f)-\nabla_{f}\log p(f)$, $\mathbf{D}(f)=(\nabla_{\mathbf{w}}f)^{2}$, $\mathbf{Q}(f)=\mathbf{0}$, $\Gamma(f)=H_{\mathbf{w}}f$. Therefore, the Fokker-Planck equation of the probability density $p(f,t)$ for $f_{t}(\cdot;\mathbf{w})$ then can be derived as
$\frac{\partial p(f,t)}{\partial t}=\nabla^{T}\cdot\left([\mathbf{D}(f)+\mathbf{Q}(f)][p(f,t)\nabla_{f}U(f)+\nabla p(f,t)]\right),$ (19)
from which the stationary distribution of $p(f,t)$ can be verified as $\exp(-U(f))=p(f|\mathcal{D})$, that is, the target posterior distribution over functions.
For the proof of Proposition 3.2, the functional Hamiltonian dynamics defined in Equation 10 can be written as
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&0\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}U(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t$ (20)
$=-\mathbf{Q}(f,g)\nabla H(f_{t},g_{t})\mathrm{d}t,$
where $\mathbf{Q}(f,g)$ is s a skew-symmetric curl matrix, $\mathbf{D}(f,g)=\mathbf{0}$. Therefore, the stationary distributions of the functional Hamiltonian dynamics defined in Equation 10 is $\pi(f,g)\propto\exp(H(f,g))$, and simply (marginalize) discard the auxiliary variables can obtain the target true posterior $p(f|\mathcal{D})$.
For the stochastic gradient version, the functional stochastic gradient Hamiltonian dynamics (fSGHMC) for the discretization update rule for samples of $\mathbf{w}$ and $\mathbf{z}$ defined in Equation 12 is defined as follows:
$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}\tilde{U}(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t+\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g)^{T}\end{array}\right]\mathrm{d}t,$ (21)
where $\mathbf{D}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$ and $\mathbf{Q}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ is the same as that in full-batch functional Hamiltonian dynamics. Then, the (marginal) stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$.
Moreover, the comparisons of matrix $\mathbf{D}(\cdot)$ and $\mathbf{Q}(\cdot)$ for naive SGMCMC methods and our functional SGMCMC are shown in Table 5.
Table 5: General framework of naive SGMCMC and functional SGMCMC algorithms.
Dynamics $X$ $G(X)$ $\mathbf{Q}(X)$ $\mathbf{D}(X)$
SGLD $\mathbf{w}$ $\tilde{U}(\mathbf{w})$ $\mathbf{0}$ $\mathbf{I}_{d}$
fSGLD $f$ $\tilde{U}(f)$ $\mathbf{0}$ $(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)$
SGHMC $(\mathbf{w},\mathbf{z})$ $\tilde{H}(\mathbf{w},\mathbf{z})$ $\left(\begin{array}[]{cc}\mathbf{0}&-\mathbf{I}_{d}\\ \mathbf{I}_{d}&\mathbf{0}\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C\end{array}\right)$
fSGHMC $(f,g)$ $\tilde{H}(f,g)$   $\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$
Appendix C Pseudocode for Functional SGMCMC.
Algorithm 1 presents the pseudocode for fSGLD. And the pseudocode for fSGHMC is shown in Algorithm 2.
Algorithm 1 Functional SGLD (fSGLD)
  Input: Dataset $\mathcal{D}$, pre-trained GP prior $p(f)$, initialized $\mathbf{w}_{0}\sim\mathcal{N}(0,I)$, step size $\epsilon_{t}$, noise $\eta_{t}\sim\mathcal{N}(0,1)$, number of burn-in iterations $N_{b}$, number of sample $K$
  ##\## Burn-in stage :
  for $t=0$ to $N_{b}$ do
     draw measurement set $\mathbf{X}_{\mathcal{M}}$;
     update $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}\tilde{U}(f_{t}))+\sqrt{2\epsilon_{t}}\eta_{t}$ using Equation 9;
  end for
  ##\## Sampling stage :
  $S\leftarrow\varnothing$;
  $\mathbf{w}_{0}\leftarrow\mathbf{w}_{N_{b}}$;
  for $t=0$ to $K$ do
     draw measurement set $\mathbf{X}_{\mathcal{M}}$;
     $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}\tilde{U}(f_{t}))+\sqrt{2\epsilon_{t}}\eta_{t}$ using Equation 9;
     $S\leftarrow S\cup\left\{\mathbf{w}_{t+1}\right\}$;
  end for
Algorithm 2 Functional SGHMC (fSGHMC)
  Input: Dataset $\mathcal{D}$, pre-trained GP prior $p(f)$, initialized $\mathbf{w}_{0}\sim\mathcal{N}(0,I)$, $\mathbf{z}_{0}\sim\mathcal{N}(0,M)$, step size $\epsilon_{t}$, noise $\eta_{t}\sim\mathcal{N}(0,1)$, number of burn-in iterations $N_{b}$, leapfrog steps $m$, number of sample $K$
  ##\## Burn-in stage :
  for $t=0$ to $N_{b}$ do
     resample momentum $\mathbf{z}_{t}\sim\mathcal{N}(0,M)$
     $(\mathbf{w}_{0},\mathbf{z}_{0})\leftarrow(\mathbf{w}_{t},\mathbf{z}_{t})$;
     for $i=1$ to $m$ do
        draw measurement set $\mathbf{X}_{\mathcal{M}}$;
        update $(\mathbf{w}_{i},\mathbf{z}_{i})$ using Equation 12;
     end for
     $(\mathbf{w}_{t+1},\mathbf{z}_{t+1})\leftarrow(\mathbf{w}_{m},\mathbf{z}_{m})$
  end for
  ##\## Sampling stage :
  $S\leftarrow\varnothing$;
  $\mathbf{w}_{0}\leftarrow\mathbf{w}_{N_{b}}$;
  for $t=0$ to $K$ do
     resample momentum $\mathbf{z}_{t}\sim\mathcal{N}(0,M)$
     $(\mathbf{w}_{0},\mathbf{z}_{0})\leftarrow(\mathbf{w}_{t},\mathbf{z}_{t})$;
     for $i=1$ to $m$ do
        draw measurement set $\mathbf{X}_{\mathcal{M}}$;
        update $(\mathbf{w}_{i},\mathbf{z}_{i})$ using Equation 12;
     end for
     $(\mathbf{w}_{t+1},\mathbf{z}_{t+1})\leftarrow(\mathbf{w}_{m},\mathbf{z}_{m})$
     $S\leftarrow S\cup\left\{\mathbf{w}_{t+1}\right\}$;
  end for
Appendix D Ablation Study
D.1 The effects of the sample size for naive SGMCMC and functional SGMCMC
In this section, we investigate the effect of sample size on the results of naive SGMCMC and our functional SGMCMC by varying the number of samples. For the 1-D extrapolation example in Section 5.1. We consider three different sample sizes for SGLD, SGHMC, fSGLD, and fSGHMC: 10, 80, and 200, respectively. The extrapolation results are shown in Figure 3. We can see that no matter the sample size, there is almost no difference in the fitting effect and uncertainty estimation for both naive SGMCMC and our functional SGMCMC, which indicates that the sample size has almost no effect on our experimental results.
(a) fSGLD-10
(b) fSGLD-80
(c) fSGLD-200
(d) SGLD-10
(e) SGLD-80
(f) SGLD-200
(g) fSGHMC-10
(h) fSGHMC-80
(i) fSGHMC-200
(j) SGHMC-10
(k) SGHMC-80
(l) SGHMC-200
Figure 3: The effect of sample size on 1-D extrapolation example for fSGLD, SGLD, fSGHMC, SGHMC. The number after the short line in each subheading represents the sample size.
For the UCI regressions in Section 5.2, the sample size is 15 for all naive SGMCMC and functional SGMCMC methods. We now consider a bigger sample size of 150 for all methods on the Yacht and Boston datasets. The RMSE and NLL results for the naive SGMCMC and functional SGMCMC are shown in Table 6 and Table 7, respectively. We randomly split each dataset into 90% training data and 10% test data, which is repeated 5 times to ensure validity. These two results have only very slight fluctuations compared to the results in Table 2 and Table 3, which still illustrates that our results are almost independent of the sample size. Our functional fSGLD and fSGHMC consistently demonstrate superior performance.
Table 6: The table shows the average RMSE results for sample size effects (150 samples).
RMSE
SGLD fSGLD SGHMC fSGHMC
Yacht 1.095 $\pm$ 0.128 0.409 $\pm$ 0.112 1.187 $\pm$ 0.141 0.246 $\pm$ 0.142
Boston 1.248 $\pm$ 0.072 0.392 $\pm$ 0.112 1.332 $\pm$ 0.068 0.270 $\pm$ 0.135
Table 7: The table shows the average NLL results for sample size effects (150 samples).
NLL
SGLD fSGLD SGHMC fSGHMC
Yacht -0.352 $\pm$ 0.073 -2.332 $\pm$ 0.493 -2.300 $\pm$ 0.181 -3.406 $\pm$ 0.517
Boston -0.509 $\pm$ 0.129 -2.068 $\pm$ 0.159 -2.352 $\pm$ 0.100 -3.060 $\pm$ 0.284
D.2 The effects of the number of measurement points on the gradient estimation of functional prior
In the 1-D extrapolation example in Section 5.1, we sampled 40 measurement points from 20 training data and an additional 40 inducing points from Uniform $(-1,1)$ to estimate the gradient of log functional prior for fSGLD and fSGHMC. We now investigate the effects of the number of measurement points on the approximate estimation of the gradient of the log functional prior in functional SGMCMC. We consider two other cases for the number of measurement points: 10 measurement points sampled from 20 training data and an additional 40 inducing points from Uniform $(-1,1)$; and 400 measurement points sampled from 20 training data and an additional 1000 inducing points from Uniform $(-1,1)$. The comparison results are shown in Figure 4. We can see that there is almost no difference in the prediction performance for the three different numbers of measurement points. The uncertainty estimations also fluctuate only slightly. Overall, we can conclude that the number of measurement points for gradient estimation of functional prior does not significantly impact the results of our methods on 1-D extrapolation example.
(a) fSGLD-m10
(b) fSGLD-m40
(c) fSGLD-m400
(d) fSGHMC-m10
(e) fSGHMC-m40
(f) fSGHMC-m400
Figure 4: The effects of the number of measurement points on the gradient estimation of functional prior for fSGLD and fSGHMC. The number after the short line in each subheading represents the number of measurement points.
In the UCI regressions in Section 5.2, for datasets with sample sizes less than 1000, we estimate the gradient of the functional prior for fSGLD and fSGHMC using the training data as the finite measurement points. For datasets with sample sizes larger than 1000, we randomly sample 1000 data from the training data as measurement points. We now consider using only 40 measurement points on the Yacht and Boston datasets. The results of RMSE and NLL are shown in Table 8 and Table 9, respectively. Compared to the results in Table 2 and Table 3, the RMSE results are slightly worse for both fSGLD and fSGHMC. The NLL results for fSGLD get slightly worse on the Yacht dataset and remain essentially unchanged on the Boston dataset. The NLL results for fSGHMC are slightly worse on both datasets but still significantly outperforms all other methods. Overall, for the UCI regressions, the more measurement points we used, the better the performance of our methods, probably because more measurement points are more accurate for the gradient estimation of the functional prior.
Table 8: The table shows the average RMSE results for the effects of the number of measurement points.
RMSE
fSGLD fSGHMC
Yacht 0.496 $\pm$ 0.072 0.297 $\pm$ 0.111
Boston 0.448 $\pm$ 0.098 0.290 $\pm$ 0.126
Table 9: The table shows the average NLL results for the effects of the number of measurement points.
NLL
fSGLD fSGHMC
Yacht -2.327 $\pm$ 0.564 -2.738 $\pm$ 0.428
Boston -2.277 $\pm$ 0.089 -2.607 $\pm$ 0.329
D.3 The effects of the kernel function on the functional GP prior
(a) fSGLD-RBF
(b) fSGLD-Linear
(c) fSGLD-Matern
(d) fSGHMC-RBF
(e) fSGHMC-Linear
(f) fSGHMC-Matern
Figure 5: The effects of the kernel function on the functional GP prior. The text after the short line in each subheading represents the corresponding kernel functions.
Functional Gaussian processes (GPs) priors are able to encode prior knowledge about function properties (e.g., periodicity and smoothness) through corresponding kernel functions. We used RBF kernel in the 1-D extrapolation experiment, which is suited for modeling polynomial functions. We now consider two other kernel functions of GP prior to our functional SGMCMC: Matern kernel and Linear kernel (unsuitable for modeling polynomial oscillatory curves). The results are shown in Figure 5. The results from the Linear kernel in Figure 5(b) and Figure 5(e) for fSGLD and fSGHMC respectively show that the fitting performance deteriorates significantly in the observation region and also demonstrate a strong linear trend (converging to a horizontal line) in the middle unseen region $[-0.25,0.25]$. The uncertainty intervals in the leftmost $[-1,-0.75]$ and rightmost $[0.75,1]$ non-observation regions deviate far from the true function. On the contrary, the results of the Matern kernel are hardly different from those of the RBF kernel, due to the fact that it also has a strong ability to describe polynomial functions.

### v2

@@
 Abstract
-Classical variational inference for Bayesian neural networks (BNNs) in parameter space usually suffers from unresolved prior issues such as knowledge encoding intractability and pathological behaviors in deep networks, which could lead to an improper posterior inference.
-Hence, functional variational inference has been proposed recently to resolve these issues via stochastic process priors.
-Beyond variational inference, stochastic gradient Markov Chain Monte Carlo (SGMCMC) is another scalable and effective inference method for BNNs to asymptotically generate samples from true posterior by simulating a continuous dynamic.
-However, the existing SGMCMC methods only work in parametric space, which has the same issues of parameter-space variational inference, and extending the parameter-space dynamics to function-space dynamics is not a trivial undertaking.
-In this paper, we introduce a new functional SGMCMC scheme via newly designed diffusion dynamics, which can incorporate more informative functional priors.
-Moreover, we prove that the stationary distribution of these functional dynamics is the target posterior distribution over functions.
-We demonstrate better performance in both accuracy and uncertainty quantification of our functional SGMCMC on several tasks compared with naive SGMCMC and functional variational inference methods.
-Machine Learning, ICML
+Classical parameter-space Bayesian inference for Bayesian neural networks (BNNs) suffers from several unresolved prior issues, such as knowledge encoding intractability and pathological behaviours in deep networks, which can lead to improper posterior inference.
+To address these issues, functional Bayesian inference has recently been proposed leveraging functional priors, such as the emerging functional variational inference.
+In addition to variational methods, stochastic gradient Markov Chain Monte Carlo (MCMC) is another scalable and effective inference method for BNNs to asymptotically generate samples from the true posterior by simulating continuous dynamics.
+However, existing MCMC methods perform solely in parameter space and inherit the unresolved prior issues, while extending these dynamics to function space is a non-trivial undertaking.
+In this paper, we introduce novel functional MCMC schemes, including stochastic gradient versions, based on newly designed diffusion dynamics that can incorporate more informative functional priors.
+Moreover, we prove that the stationary measure of these functional dynamics is the target posterior over functions.
+Our functional MCMC schemes demonstrate improved performance in both predictive accuracy and uncertainty quantification on several tasks compared to naive parameter-space MCMC and functional variational inference.
 1 Introduction
-Compared to regular deep neural networks that rely on a single pattern of model parameters only, Bayesian neural networks (BNNs) have demonstrated the ability to provide a principled predictive uncertainty estimation and enhanced generalization performance through a model integration under posterior distribution over network weights (Neal, 1995; Gal et al., 2016; Wilson & Izmailov, 2020).
-With the continued development and success of Bayesian inference methods in the deep learning community, BNNs have been applied in a wide variety of domains and tasks, including prediction in astrophysics (Cranmer et al., 2021), medical diagnosis and diabetic detection (Filos et al., 2019; Band et al., 2022), and other safety-critical environments (Rudner et al., 2022).
-Prior choosing is apparently crucial for BNNs.
-The existing common practice is i.i.d.
-Gaussian priors for all network weights, which is also named parameter-space prior.
-However, this parameter-space prior has been shown to have some problematic issues.
-For example, as network depth increases, the function samples of Gaussian prior over weights tend to be horizontal (Duvenaud et al., 2014; Matthews et al., 2018), and the prior distribution of unit outputs becomes more heavy-tailed with ReLU activation (Vladimirova et al., 2019; Tran et al., 2022), which will lead to improper posterior inference (Papamarkou et al., 2022).
-Moreover, the effects of this parameter space prior to the posterior inference remain an open question due to the complex network architecture (Fortuin et al., 2021; Wild et al., 2022), so it is difficult to transform the valid prior knowledge into corresponding network weights’ prior.
-Given these unresolved prior problems for parameters, Bayesian inference directly in function space has gained increasing interest in recent years since one can incorporate more informative stochastic processes as the prior.
-For example, Gaussian processes (GPs) which can easily encode the prior knowledge about the potential functions via kernel functions choosing have been widely used as the prior in function-space variational inference for BNNs (Sun et al., 2019; Ma & Hernández-Lobato, 2021; Rudner et al., 2022; Pielok et al., 2022) which has been shown to yield better prediction and improved uncertainty estimation performance than variational inference in parameter space.
-For the posterior inference given prior, Markov Chain Monte Carlo (MCMC) is another valid method for BNNs apart from the commonly used variational inference (Blundell et al., 2015; Sun et al., 2019).
-Instead of optimizing a variational posterior by minimizing the distance (divergence) with the true posterior (Graves, 2011; Blundell et al., 2015; Gal & Ghahramani, 2016), MCMC approaches are regarded as the gold standard for Bayesian inference since they generate discrete samples from a continuous-time Markov chain whose stationary distribution is exactly the true posterior (Neal et al., 2011; Welling & Teh, 2011; Li et al., 2016).
-So, they do not need to make strong assumptions about the distributional form of the posterior (Alexos et al., 2022; Cobb & Jalaian, 2021).
-Even better, the computationally expensive MCMC methods became feasible for modern BNNs with large data following the introduction of their scalable stochastic gradient variants (SGMCMC) (Welling & Teh, 2011; Patterson & Teh, 2013; Chen et al., 2014; Zhang et al., 2020; Nemeth & Fearnhead, 2021).
-However, all existing MCMC methods for BNNs are still in parameter space, and extending them to function space is not a minor process because we need to design novel dynamics in function space and ensure the new dynamics would have the target posterior as the stationary distribution.
-In this work, we propose a functional SGMCMC scheme based on the fundamental Itô lemma for stochastic calculus (Itô, 1951; Brzeźniak et al., 2008; Øksendal, 2013).
-The dynamics of BNNs in function space are designed by proposing a new potential energy function where we integrate the more informative functional prior, such as GPs, into the diffusion process.
-Furthermore, we prove that the stationary distribution of the new functional dynamics is the desired posterior distribution over functions.
+Compared to traditional deep neural networks that rely on a single pattern of model parameters only, Bayesian neural networks (BNNs) demonstrate principled predictive uncertainty estimation and improved generalization by integrating models under the posterior distribution over network parameters (Neal,, 1995; Gal et al.,, 2016; Wilson and Izmailov,, 2020).
+As Bayesian inference methods continue to advance and gain success within the deep learning community, BNNs have been applied in a wide variety of domains and tasks, including dissolution prediction of planetary systems (Cranmer et al.,, 2021), medical diagnosis as diabetic detection (Filos et al.,, 2019; Band et al.,, 2022), classification of radio galaxies (Mohan and Scaife,, 2024), and other safety-critical environments (Rudner et al.,, 2022).
+Choosing an appropriate prior is critical for BNNs.
+The common practice is to use isotropy Gaussian priors for all network parameters, often referred to as parameter-space priors.
+However, these parameter-space priors have shown several problematic issues.
+For instance, as the depth of the network increases, the function samples of Gaussian prior over parameters tend to be horizontal (Duvenaud et al.,, 2014; Matthews et al.,, 2018).
+Additionally, with ReLU activations, the prior distribution of unit outputs becomes more heavy-tailed for deeper architectures.(Vladimirova et al.,, 2019; Tran et al.,, 2022).
+Furthermore, the effects of this parameter-space prior to posterior inference remain an open question due to the non-interpretability of network parameters and the complexity of network architectures Fortuin et al., (2021); Wild et al., (2022), which make it challenging to translate valid prior knowledge into a corresponding prior distribution over network parameters.
+Due to these unresolved issues with parameter-space priors, there has been growing interest in performing Bayesian inference directly in function space, where more informative stochastic processes can serve as priors.
+For example, Gaussian processes (GPs), which can easily encode prior knowledge about the potential functions via kernel functions, have been widely used as priors in functional variational inference for BNNs (Sun et al.,, 2019; Ma and Hernández-Lobato,, 2021; Rudner et al.,, 2022; Pielok et al.,, 2022).
+These functional variational inference approaches have shown improved performance compared to parameter-space variational inference.
+For posterior inference given a prior, Markov Chain Monte Carlo (MCMC) is another valid method for BNNs apart from variational inference.
+Unlike variational inference, which makes strong assumptions about the posterior distribution, non-parametric MCMC methods are asymptotically exact, making them the gold standard for Bayesian inference (Alexos et al.,, 2022).
+Even better, the computationally expensive MCMC methods became feasible for modern BNNs with large data following the introduction of their scalable stochastic gradient variants (SGMCMC) (Welling and Teh,, 2011; Patterson and Teh,, 2013; Chen et al.,, 2014; Zhang et al.,, 2020; Nemeth and Fearnhead,, 2021).
+However, existing MCMC methods for BNNs typically perform in parameter space and are based on the diffusion dynamics designed in terms of the posterior over parameters, inheriting similar prior issues as parameter-space variational inference.
+Extending these methods to function space, however, is a non-trivial task.
+It requires the design of novel dynamics that not only sample effectively from the function space but also guarantee that the resulting stationary distribution corresponds to the target posterior over functions.
+This involves addressing the inherent complexities of functional spaces while ensuring the correctness and feasibility of the sampling process.
+In this work, we propose novel functional MCMC schemes grounded in the fundamental Itô lemma for stochastic calculus (Itô,, 1951; Brzeźniak et al.,, 2008; Øksendal,, 2013).
+Specifically, the dynamics of BNNs in function space are designed by introducing new potential energy functionals that can incorporate functional priors, such as GPs, into the diffusion processes and simultaneously guarantee the target posterior over functions as the stationary measure.
 Our main contributions are as follows:
-  • We propose a new functional SGMCMC approach for BNNs that can utilize the more informative functional prior to the posterior inference.
-  • We prove that under our designed functional SGMCMC dynamics, the stationary distribution would be the target posterior distribution over functions of BNNs.
-  • Our method shows improved predictive performance and uncertainty estimation in several benchmark tasks compared to parameter-space SGMCMC and parameter- and function-space variational inference methods.
+  • We propose novel functional diffusion dynamics, along with their stochastic gradient versions for BNNs that effectively incorporate more informative functional priors into posterior inference.
+Additionally, we derive corresponding tractable MCMC samplers in parameter space, making them computationally feasible while maintaining the expressiveness of functional priors.
+  • We prove that the stationary measure would be the target posterior over functions under our designed functional Langevin dynamics and functional Hamiltonian dynamics.
+  • Our methods demonstrate improved predictive performance and uncertainty estimation in several benchmark tasks compared to parameter-space SGMCMC and parameter-/function-space variational inference methods.
 2 Preliminaries
 2.1 Bayesian Neural Networks
-[…]
+Bayesian neural networks (BNNs) place probability distribution over network parameters rather than single fixed values used in standard neural networks.
+Given a dataset $\mathcal{D}=\left\{x_{i},y_{i}\right\}_{i=1}^{N}=\left\{\mathbf{X_{\mathcal{D}}},\mathbf{Y_{\mathcal{D}}}\right\}$ , where $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ are the training inputs and $y_{i}\in\mathcal{Y}\subseteq\mathbb{R}^{c}$ are the corresponding targets, BNNs are stochastic neural networks characterized by random network parameters $\mathbf{w}\in\mathbb{R}^{k}$, treated as a k-dimensional multivariate random variable defined on a probability space $(\Omega,\mathcal{A},P)$.
+The prior distribution over parameters is denoted by $p_{0}(\mathbf{w})$, and the likelihood evaluated on the training data is represented by the conditional distribution $p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})$.
+The posterior over parameters can then be inferred using Bayes’ theorem, yielding the canonical form $p(\mathbf{w}|\mathcal{D})\propto p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})p_{0}(\mathbf{w})$.
+Furthermore, the predictive distribution for test data $\left\{x^{*},y^{*}\right\}$ is given by the expectation under the posterior as $p(y^{*}|x^{*},\mathcal{D})=\int p(y^{*}|x^{*},\mathbf{w})p(\mathbf{w}|\mathcal{D})\mathrm{d}\mathbf{w}$.
+All notations used in this paper are listed in Table 5 in Appendix A.
 2.2 Stochastic Gradient MCMC for BNNs
-Traditional MCMC methods, such as the random-walk Metropolis-Hastings (MH) algorithm (Metropolis et al., 1953), are easy to apply but suffer from a slow mixing rate and high correlation between samples.
-In contrast, the dynamics-based sampling methods utilizing gradients of log posterior density can provide a more efficient exploration of the posterior (Girolami & Calderhead, 2011).
-Also, their stochastic gradient variants (Welling & Teh, 2011; Chen et al., 2014; Baker et al., 2019), such as the stochastic gradient Langevin dynamics (SGLD) and the stochastic gradient Hamiltonian Monte Carlo (SGHMC), can scale to large datasets for practical applications.
+Dynamics-based sampling methods that leverage the gradients of the target log posterior can offer an efficient exploration of the parameter space (Ma et al.,, 2015; Izmailov et al.,, 2021; Li and Zhang,, 2024).
+Furthermore, their stochastic gradient variants, such as the stochastic gradient Langevin dynamics (SGLD) (Welling and Teh,, 2011) and the stochastic gradient Hamiltonian Monte Carlo (SGHMC) (Chen et al.,, 2014; Baker et al.,, 2019), are capable of scaling to large datasets for practical applications.
 SGLD.
-The objective of MCMC is to obtain random samples of the posterior distribution of model parameters (with the unnormalized form) $p(\mathbf{w}|\mathcal{D})\propto\exp(-U(\mathbf{w}))$, where $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})$ is the potential energy function.
-The idea of dynamics-based sampling for $p(\mathbf{w}|\mathcal{D})$ is to simulate a continuous-time stochastic process whose stationary distribution is the target posterior.
-Combining Langevin dynamics (Neal et al., 2011) with stochastic Robbins-Monro optimization algorithm (Robbins & Monro, 1951), Welling & Teh (2011) proposed SGLD for scalable MCMC algorithms.
+The objective of MCMC for BNNs is to generate random samples from the posterior distribution over parameters (with the unnormalized form) as $p(\mathbf{w}|\mathcal{D})\propto\exp(-U(\mathbf{w}))$, where $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})-\log p_{0}(\mathbf{w})$ is the potential energy function.
+By combining Langevin dynamics (Neal et al.,, 2011) with the stochastic Robbins-Monro optimization algorithm (Robbins and Monro,, 1951), Welling and Teh, (2011) proposed SGLD as a scalable MCMC algorithm.
 The approximated discretization transition rule for posterior samples is
 $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})+\sqrt{2\epsilon_{t}}\eta_{t},$ (1)
-where $\epsilon_{t}$ is the step-size, $\eta_{t}$ is a standard Gaussian, and $\nabla\tilde{U}(\mathbf{w})$ uses mini-batch stochastic gradient for the likelihood term as $\nabla\tilde{U}(\mathbf{w})=-\frac{N}{n}\sum_{i=1}^{n}\nabla\log p(y_{i}|x_{i},\mathbf{w})-\nabla\log p(\mathbf{w})$ and $n\ll N$.
-With the decreasing step-size $\epsilon_{t}$, the stochastic gradient noise tends to zero faster than the injected Gaussian noise $\eta_{t}$, and the discretization error is eliminated, then the MH step can be ignored.
+where $\epsilon_{t}$ represents the step-size, and $\eta_{t}$ is a standard Gaussian noise term.
+The gradient $\nabla\tilde{U}(\mathbf{w})$ is computed using a mini-batch stochastic gradient for the likelihood term given by $\nabla\tilde{U}(\mathbf{w})=-\frac{N}{n}\sum_{i=1}^{n}\nabla\log p(y_{i}|x_{i};\mathbf{w})-\nabla\log p(\mathbf{w})$ with $n\ll N$.
+As the step-size $\epsilon_{t}\rightarrow 0$, the stochastic gradient noise vanishes more quickly than the injected Gaussian noise $\eta_{t}$, which ensures the discretization error is eliminated.
+Hence, the Metropolis-Hastings (MH) correction step can be ignored.
 SGHMC.
-To enable a more efficient exploration in the state space, Hamiltonian dynamics (Duane et al., 1987; Neal, 2012) introduces a momentum variable $\mathbf{z}\in\mathbb{R}^{d}$, and the new joint distribution is $p(\mathbf{w}|\mathcal{D})p(\mathbf{z})\propto\exp(-H(\mathbf{w},\mathbf{z}))$ with corresponding energy function $H(\mathbf{w},\mathbf{z})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})-\log p(\mathbf{z})$, where $p(\mathbf{z})$ is commonly assumed as a zero-mean Gaussian $\mathcal{N}(\mathbf{z}|0,M)$.
-The sample rule in the full-batch Hamiltonian Monte Carlo is defined as following with an MH step (Neal, 2012; Cobb & Jalaian, 2021):
+To enable more efficient exploration of the state space, Hamiltonian dynamics (Duane et al.,, 1987; Neal,, 2012) introduces a momentum variable $\mathbf{z}\in\mathbb{R}^{k}$.
+The new target joint distribution is $p(\mathbf{w}|\mathcal{D})p(\mathbf{z})\propto\exp(-H(\mathbf{w},\mathbf{z}))$ with corresponding energy function $H(\mathbf{w},\mathbf{z})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}},\mathbf{w})-\log p(\mathbf{w})-\log p(\mathbf{z})$, where $p(\mathbf{z})$ is commonly assumed to be a zero-mean Gaussian $\mathcal{N}(\mathbf{z}|0,M)$.
+The sampling rule in the full-batch Hamiltonian Monte Carlo is defined as following with an MH step (Neal,, 2012; Cobb and Jalaian,, 2021):
 $\mathbf{w}_{t+1}$ $=\mathbf{w}_{t}+\epsilon_{t}M^{-1}\mathbf{z}_{t},$ (2)
 $\mathbf{z}_{t+1}$ $=\mathbf{z}_{t}-\epsilon_{t}\nabla U\left(\mathbf{w}_{t}\right).$
-Like SGLD, SGHMC uses stochastic gradient $\nabla\tilde{U}(\mathbf{w})$ for the sampling of momentum $\mathbf{z}$, also with an additional friction term $\epsilon_{t}M^{-1}\mathbf{z}_{t}$ and a Gaussian injected noise to guarantee the (marginal) stationary distribution of the Hamiltonian dynamics still be the target posterior distribution (Chen et al., 2014).
-The specific update rule for samples of $\mathbf{z}$ is as: $\mathbf{z}_{t+1}=\mathbf{z}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})-\epsilon_{t}CM^{-1}\mathbf{z}_{t}+\sqrt{2C\epsilon_{t}}\eta_{t}$, where $C>0$ is a scalar coefficient.
-3 Functional Stochastic Gradient MCMC
-Generally, dynamics-based MCMC approaches are derived from the Itô diffusion (Øksendal, 2003), which is similar to the Brownian motion of the particle subjected to a potential in fluid physics.
-In this section, to solve the parameter-space prior issues in existing SGMCMC methods, we propose a new functional SGMCMC scheme via a newly designed diffusion dynamics, which is capable of incorporating more informative functional priors.
-In Section 3.1, we first review the Itô diffusion for dynamics-based MCMC.
-Then we design a new functional Langevin dynamics directly for the function mapping defined by the BNNs.
-We prove that the stationary distribution of such functional Langevin dynamics is the target posterior distribution over functions.
-Furthermore, in Section 3.2, we design a functional Hamiltonian dynamics and the corresponding stochastic gradient sampling scheme for BNNs.
-We also prove it can guarantee the desired invariant distribution.
-3.1 Functional SGLD for BNNs
-An Itô diffusion in n-dimensional Euclidean space is defined as a stochastic differential equation (SDE) to describe the evolution of a continuous-time stochastic process $X:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}$ on a probability space $(\Omega,\Sigma,P)$.
-The general form of an Itô diffusion driven by the standard Wiener process (Brownian motion) is as follows:
-$\mathrm{d}X_{t}=\mu(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},$ (3)
-where $X_{t}$ represents the state of the stochastic process at time $t$, $B\in\mathbb{R}^{n}$ is an n-dimensional Brownian motion, $\mu(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a vector field denoting the deterministic drift term and matrix field $\sigma(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ is known as the diffusion coefficient for $X_{t}$.
-Both $\mu(\cdot)$ and $\sigma(\cdot)$ are assumed to satisfy the usual Lipschitz continuity condition (Ghosh, 2010).
-And $\mathrm{d}B_{t}$ is the increment of a Wiener process distributed as $\mathrm{d}B_{t}\sim\mathcal{N}(0,\mathrm{d}t\cdot\mathbf{I}_{n})$.
-Let $p(x,t)$ be the probability density of random variable $X_{t}$, and the corresponding evolution of this probability density of the diffusion process can be described by the Fokker–Planck equation (Risken, 1996), which is the key to derive the dynamics-based MCMC (see Appendix B).
-Naturally, the idea of existing dynamics-based SGMCMC methods for BNNs is to simulate a continuous diffusion of network weight whose stationary distribution is the target posterior $p(\mathbf{w}|\mathcal{D})$ with appropriately designed drift term and diffusion coefficient.
-To incorporate informative functional prior into the posterior inference process, we now directly consider the dynamics of function mapping $f(\cdot;\mathbf{w})$ defined by BNNs.
-According to Itô Lemma (Itô, 1951), $f(\cdot;\mathbf{w})$ is itself an Itô diffusion (Brzeźniak et al., 2008) (see Appendix A).
-That is, for exiting SGMCMC methods, the stochastic process $X$ simulated in Itô diffusion is usually for network weight or its augmentation, which we now lift to the stochastic function mapping of BNNs directly as
-$\mathrm{d}f_{t}=\mu(f_{t})\mathrm{d}t+\sigma(f_{t})\mathrm{d}B_{t}.$ (4)
-In a probability space $(\Omega,\mathcal{F},P)$ (separable and complete Polish space), suppose $p(f)$ is the prior distribution over functions, such as the classical Gaussian process (GP) prior that can easily encode prior knowledge about function properties.
-$p(\mathbf{Y_{\mathcal{D}}}|f(\mathbf{X_{\mathcal{D}}}))$ is the likelihood function with observation $\mathcal{D}=\left\{\mathbf{X_{\mathcal{D}}},\mathbf{Y_{\mathcal{D}}}\right\}$ and $p(f|\mathcal{D})$ is the posterior over functions.
-For brevity, we abbreviate them as $p(f)$, $p(\mathcal{D}|f)$ and $p(f|\mathcal{D})$, respectively.
-Let $U(f)=-\log p(\mathcal{D}|f)-\log\frac{\mathrm{d}p}{\mathrm{d}r_{\mathcal{H,F}}}(f))$ be the potential energy functional, where $r_{\mathcal{H,F}}$ is the corresponding Hausdorff / Riemannian measure, and $\mu(f)=-\nabla_{f}U(f),\sigma(f)=\sqrt{2}\cdot\mathbf{I}_{\infty}$ for the function mapping $f$ of BNNs, we have the following functional Langevin dynamics:
-$\mathrm{d}f_{t}=-\nabla_{f}U(f_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$ (5)
-and it is easy to verify that the stationary distribution for $p(f,t)$ is exactly the target posterior distribution over functions as $\pi(f)=\exp(-U(f))=p(f|\mathcal{D})$.
-However, even for the finite input case, it is intractable to discretize and sample the above diffusion of function mapping in practice directly.
-Considering that a function mapping $f(\cdot;\mathbf{w})$ of BNNs is completely determined by the network weights $\mathbf{w}$ (assuming the network architecture is fixed), we solve this issue by transforming the functional Langevin dynamics to a weight-space dynamics of which stationary distribution is still $p(f|\mathcal{D})$.
-We firstly design the following specific Riemannian functional Langevin dynamics for $f(\cdot;\mathbf{w})$ as follows:
-$\mathrm{d}f_{t}(\cdot;\mathbf{w})=$ $\mu(f_{t}(\cdot;\mathbf{w}))\mathrm{d}t+\sigma(f_{t}(\cdot;\mathbf{w}))\mathrm{d}B_{t}$ (6)
-$=$ $\left[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f\right]\mathrm{d}t$
-$+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t},$
-where $\nabla_{\mathbf{w}}f$ is the gradient of $f(\cdot;\mathbf{w})$ w.r.t.
-weight $\mathbf{w}$, $H_{\mathbf{w}}f$ is the Hessian matrix of $f(\cdot;\mathbf{w})$ w.r.t.
-$\mathbf{w}$.
-The stationary distribution of this dynamics is guaranteed to be $p(f|\mathcal{D})$ by the following Proposition.
+Like SGLD, SGHMC uses stochastic gradient $\nabla\tilde{U}(\mathbf{w})$ for the sampling of momentum $\mathbf{z}$, but with an additional friction term $\epsilon_{t}M^{-1}\mathbf{z}_{t}$ and a Gaussian injected noise to ensure that the (marginal) stationary distribution of the Hamiltonian dynamics remains the target posterior distribution (Chen et al.,, 2014).
+The specific update rule for the momentum $\mathbf{z}$ is given by $\mathbf{z}_{t+1}=\mathbf{z}_{t}-\epsilon_{t}\nabla\tilde{U}(\mathbf{w}_{t})-\epsilon_{t}CM^{-1}\mathbf{z}_{t}+\sqrt{2C\epsilon_{t}}\eta_{t}$, where $C>0$ is a scalar coefficient.
+3 Functional Dynamics-based MCMC
+In this section, we address the prior issues present in parameter-space MCMC by proposing two novel functional MCMC schemes based on newly designed diffusion dynamics, which allow for the incorporation of more informative functional priors.
+In Section 3.1, we introduce novel functional Langevin dynamics designed directly for the random function mapping defined by BNNs, along with a corresponding tractable stochastic gradient MCMC sampler for network parameters.
+We demonstrate that the stationary probability measure of this functional Langevin dynamics corresponds to the target posterior over functions.
+Additionally, in Section 3.2, we develop functional Hamiltonian dynamics also with a stochastic gradient version for BNNs.
+We prove that this scheme also guarantees the desired functional posterior as the stationary measure.
+3.1 Functional Langevin Dynamics for BNNs
+Dynamics-based MCMC techniques are grounded in the general framework of Itô diffusion (Øksendal,, 2003).
+The main idea behind parameter-space MCMC methods for BNNs is to simulate a continuous diffusion process of network parameters $\mathbf{w}$ such that its stationary distribution is the target posterior $p(\mathbf{w}|\mathcal{D})$, using appropriately designed drift term and diffusion coefficient (see Appendix B).
+Naturally, to incorporate more informative functional prior into the posterior inference process, we now consider lifting the parameter-space Langevin dynamics onto function space.
+Suppose $f(\cdot;\mathbf{w}):\mathcal{X}\times\mathbb{R}^{k}\rightarrow\mathcal{Y}$ is the product-measurable random function mapping defined by a BNN.
+Assums $f(\cdot;\mathbf{w})\in\mathbb{H}$, where $\mathbb{H}$ is an infinite-dimensional function (Polish) space with Borel $\sigma$-algebra $\mathcal{B}(\mathbb{H})$.
+The likelihood is defined as $p$: $\mathcal{Y}\times\mathbb{H}\rightarrow[0,\infty)$, mapping $(\mathcal{Y},f(\mathcal{X};\mathbf{w}))\mapsto p(\mathcal{Y}|f(\mathcal{X};\mathbf{w}))$, where $\mathcal{Y}\subseteq\mathbb{R}^{c}$ is Borel measurable and $p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$ denotes the likelihood evaluated on the training data.
+Let $\mathcal{P}(\mathbb{H})$ represent the space of Borel probability measures on $\mathcal{B}(\mathbb{H})$, and assume the functional prior measure $P_{0}\in\mathcal{P}(\mathbb{H})$ with topological support $\operatorname{supp}(P_{0})$, such as the classical GP prior denoted as $P_{0}\sim\mathcal{GP}(\mathbf{m},\mathbf{K})$.
+The posterior measure $P_{f|\mathcal{D}}\in\mathcal{P}(\mathbb{H})$, induced by the functional prior and likelihood, is defined by the Radon–Nikodym derivative as $P_{f|\mathcal{D}}(\mathrm{d}f)\propto\exp(-\Phi(f))P_{0}(\mathrm{d}f)$, where $\Phi(f)$ is the negative log-likelihood as $-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$ (Matthews et al.,, 2016; Lambley,, 2023).
+According to Itô Lemma (Itô,, 1951), $f(\cdot;\mathbf{w})$ itself follows an Itô diffusion (Brzeźniak et al.,, 2008) (see Appendix B).
+Thus, we propose to directly simulate the diffusion process of the random function mapping $f(\cdot;\mathbf{w})$ of BNNs, where one can incorporate informative functional prior, $P_{0}$.
+Suppose $\varnothing\neq E\subseteq\operatorname{supp}(P_{0})$, and let $I_{0}:E\rightarrow\mathbb{R}$ be an Onsager–Machlup (OM) functional for $P_{0}$, which can be heuristically interpreted as the negative logarithm of $P_{0}$ (Ayanbayev et al.,, 2021) (see Appendix B).
+Similar to parameter-space Langevin dynamics, the Langevin dynamics in function space can be generally represented by the stochastic differential equation (SDE): $\mathrm{d}f_{t}=-\nabla U(f_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}$, where $B$ is a standard Wiener process ((Brownian motion), and potential energy functional $U:E\rightarrow\mathbb{R}$ is given by $U(f)=\Phi(f)+I_{0}(f)$.
+This functional $U(f)$ is proven to be an OM functional for $P_{f|\mathcal{D}}$ (Lambley,, 2023), ensuring that the target posterior measure $P_{f|\mathcal{D}}$ as the stationary measure $\pi(f)$.
+However, discretizing and sampling from this diffusion process in function space is intractable due to the infinite-dimensional nature of the problem.
+Since $f(\cdot;\mathbf{w})$ is completely determined by the network parameters $\mathbf{w}$ (given a fixed network architecture), we address this challenge by designing a specific functional Langevin dynamics for $f(\cdot;\mathbf{w})$, which maintains the stationary measure $P_{f|\mathcal{D}}$.
+We then transform this into the corresponding parameter-space dynamics for network parameters.
+Suppose that function space $\mathbb{H}$ is equipped with a Riemannian manifold structure (Wang et al.,, 2022), we first design the following functional Riemannian Langevin dynamics for $f(\cdot;\mathbf{w})$ on $\mathbb{H}$:
+$\mathrm{d}f_{t}(\cdot;\mathbf{w})=\mu(f_{t}(\cdot;\mathbf{w}))\mathrm{d}t+\sqrt{2G^{-1}(f_{t}(\cdot;\mathbf{w}))}\mathrm{d}B_{t}$ (3)
+$=\Big{[}-(\nabla_{\mathbf{w}}f_{t})^{T}(\nabla_{\mathbf{w}}f_{t})\nabla_{f}(-\log p(\mathbf{Y}_{\mathcal{D}}|f_{t}(\mathbf{X}_{\mathcal{D}};\mathbf{w}))+$
+$\quad I_{0}(f_{t}))+H_{\mathbf{w}}f_{t}\Big{]}\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f_{t})^{T}\mathrm{d}B_{t},$
+where $\nabla_{\mathbf{w}}f$ denotes the gradient of $f(\cdot;\mathbf{w})$ w.r.t.
+parameters $\mathbf{w}$, $H_{\mathbf{w}}f$ is the Hessian matrix, and $G^{-1}(f(\cdot;\mathbf{w}))=(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)$.
+The stationary probability measure of this dynamics is guaranteed to be $P_{f|\mathcal{D}}$ as stated in the following proposition.
 Proposition 3.1.
-The stationary distribution of the functional Langevin dynamics defined in Equation 6 is the target posterior distribution over functions $p(f|\mathcal{D})$.
+The stationary probability measure of the functional Langevin dynamics defined in Equation 3 is the target posterior over functions $P_{f|\mathcal{D}}$.
 Proof.
-The proof is given in Appendix B. ∎
-Secondly, we achieve the function-space to weight-space transformation based on the relationship between the dynamics of a random variable with the dynamics of its function.
-According to the Itô Lemma (Itô, 1951), given the general diffusion framework for weights as $\mathrm{d}\mathbf{w}_{t}=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$, the differential of the function mapping $f(\cdot;\mathbf{w})$ can be denoted as
-$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ (7)
-$=(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}\mathbf{w}_{t}+\frac{1}{2}(\mathrm{d}\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\mathrm{d}\mathbf{w}_{t}$
+The proof is given in Appendix C. ∎
+To establish a tractable functional MCMC scheme, we then transform the proposed functional Riemannian Langevin dynamics for $f(\cdot;\mathbf{w})$ into the corresponding Langevin dynamics for the network parameters $\mathbf{w}$.
+Note that the differential of the function mapping $f(\cdot;\mathbf{w})$ is given by the Itô Lemma (Itô,, 1951) as
+$\mathrm{d}f_{t}(\cdot;\mathbf{w})=(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}\mathbf{w}_{t}+\frac{1}{2}(\mathrm{d}\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\mathrm{d}\mathbf{w}_{t}$ (4)
 $=\left((\nabla_{\mathbf{w}}f)^{T}\mu(\mathbf{w}_{t})+\frac{1}{2}Tr[\sigma(\mathbf{w}_{t})^{T}(H_{\mathbf{w}}f)\sigma(\mathbf{w}_{t})]\right)\mathrm{d}t$
-$+(\nabla_{\mathbf{w}}f)^{T}\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}.$
-Given our defined functional dynamics in Equation 6 and the above relationship, the corresponding diffusion process of network weights in parameter space under our functional Langevin dynamics can be derived as
+$+(\nabla_{\mathbf{w}}f)^{T}\sigma(\mathbf{w}_{t})\mathrm{d}B_{t},$
+where $\mathrm{d}\mathbf{w}_{t}=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$ is the general diffusion form for $\mathbf{w}$.
+By combining our functional Langevin dynamics defined in Equation 3 with this differential relationship between $f(\cdot;\mathbf{w})$ and $\mathbf{w}$, it is straightforward to reversely derive the corresponding diffusion process for $\mathbf{w}$ in parameter spaces as follows:
 $\mathrm{d}\mathbf{w}_{t}=\mu(\mathbf{w}_{t})\mathrm{d}t+\sigma(\mathbf{w}_{t})\mathrm{d}B_{t}$ (5)
-$=\nabla_{\mathbf{w}}f[\nabla_{f}(\log p(\mathcal{D}|f_{t})+\log\frac{\mathrm{d}p_{0}}{\mathrm{d}r_{\mathcal{H,F}}}(f))]\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$
-where the corresponding drift term $\mu(\mathbf{w}_{t})=-\nabla_{\mathbf{w}}U(f_{t})$, and the diffusion coefficient $\sigma(\mathbf{w}_{t})=\sqrt{2}\cdot\mathbf{I}_{d}$.
-Then, the function samples of the target posterior $p(f|\mathcal{D})$ as the stationary distribution of the functional Langevin dynamics can be obtained by discretizing and sampling the diffusion process of weights in Equation 8.
-In practice, the stochastic gradient variant drift term of SDE in Equation 8 is denoted as $-\nabla_{\mathbf{w}}\tilde{U}(f_{t})$, which uses the minibatch version for the likelihood term.
-For the gradient of the log prior distribution over functions, we firstly assume an approximated GP prior for BNNs, since in the limit of infinite width, the prior measure of BNNs converges to a GP with NNGP kernel (Neal, 1995; Lee et al., 2017; Matthews et al., 2018; Lee et al., 2019).
-Due to the infinite-dimensional property of functional distribution, we solve the gradient of functional prior on finite measurement points $\mathbf{X}_{\mathcal{M}}\stackrel{{\scriptstyle\text{ dif }}}{{=}}\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\right]^{\mathrm{T}}$ as $\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f^{\mathbf{X}_{\mathcal{M}}})$, where $f^{\mathbf{X}_{\mathcal{M}}}$ are corresponding function values evaluated at $\mathbf{X}_{\mathcal{M}}$.
-Then the prior $p(f^{\mathbf{X}_{\mathcal{M}}})$ is a multivariate Gaussian that could be solved analytically for the gradient.
-The specific discretization update rule for weight samples under functional stochastic gradient Langevin dynamics (fSGLD) is
-$\mathbf{w}_{t+1}=$ $\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\Big{[}-\frac{N}{n}\sum_{i=1}^{n}\nabla_{f}\log p(y_{i}|f_{t}(x_{i};\mathbf{w}_{t}))-$ (9)
-$\hskip 62.59596pt\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log p(f_{t}^{\mathbf{X}_{\mathcal{M}}})\Big{]}+\sqrt{2\epsilon_{t}}\eta_{t}.$
-Similar to the naive SGLD, with the discretization step-size $\epsilon_{t}$ decays to zero, the stationary distribution for functional $p(f,t)$ of fSGLD continues to be the target posterior $p(f|\mathcal{D})$ over functions.
-Note that although the function samples of posterior $p(f|\mathcal{D})$ are drawn by the discretization process for corresponding diffusion of weights in Equation 9, our fSGLD is different from the existing SGLD in parameter space, which means the stationary distributions from two dynamics would be different as well.
-Compared to the naive parameter-space SGLD in Equation 1 controlled only by the limited parameter information, the discretization update rule for weight samples in Equation 9 under our fSGLD scheme contains information about the function mapping, which is important for accurate posterior inference.
-The pseudocode for fSGLD is shown in Algorithm 1 in Appendix C.
-3.2 Functional SGHMC for BNNs
-For the naive Hamiltonian dynamics, $X=(\mathbf{w},\mathbf{z})$ is the augmentation of model parameters.
-We now let $X=(f,g)$ and define the Hamiltonian dynamics in function space as $H(f,g)=U(f)-\log p(g)$, where $p(g)$ is the functional auxiliary distribution induced by the distribution of auxiliary variable $\mathbf{z}$.
-We design the functional Hamiltonian dynamics for $f(\cdot;\mathbf{w})$ and $g(\cdot;\mathbf{z})$ as
-$\mathrm{d}f_{t}(\cdot;\mathbf{w})=-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (10)
-$\mathrm{d}g_{t}(\cdot;\mathbf{z})=-(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f\nabla_{f}U(f_{t})\mathrm{d}t.$
+$=-\nabla_{\mathbf{w}}f_{t}\left[\nabla_{f}(-\log p(\mathbf{Y}_{\mathcal{D}}|f_{t}(\mathbf{X}_{\mathcal{D}};\mathbf{w}))+I_{0}(f_{t}))\right]\mathrm{d}t$
+$+\sqrt{2}\mathrm{d}B_{t},$
+where the drift term $\mu(\mathbf{w})=-\nabla_{\mathbf{w}}U(f)$ , and the diffusion coefficient $\sigma(\mathbf{w}_{t})=\sqrt{2}\cdot\mathbf{I}_{k}$.
+Then, by discretizing and sampling from this corresponding diffusion process for the network parameters, we can obtain function samples from the target posterior measure $P_{f|\mathcal{D}}$ in a tractable manner.
+In practice, the stochastic gradient variant for the drift term of SDE in Equation 5 is denoted as $-\nabla_{\mathbf{w}}\tilde{U}(f)$, which utilizes the minibatch computation for the likelihood term.
+To estimate the gradient of the prior measure, we assume an approximate GP prior for BNNs.
+This is motivated by the fact that, in the limit of infinite width, the prior measure of BNNs converges to a GP with the Neural Network Gaussian Process (NNGP) kernel (Neal,, 1995; Lee et al.,, 2017; Matthews et al.,, 2018; Lee et al.,, 2019).
+Due to the infinite-dimensional nature of the functional distribution, we solve the gradient of functional prior on finite measurement points $\mathbf{X}_{\mathcal{M}}\stackrel{{\scriptstyle\text{ def }}}{{=}}\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{M}\right]^{\mathrm{T}}$ as $\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log P_{0}(f^{\mathbf{X}_{\mathcal{M}}})$, where $f^{\mathbf{X}_{\mathcal{M}}}$ are corresponding function values evaluated at $\mathbf{X}_{\mathcal{M}}$.
+The prior $P_{0}(f^{\mathbf{X}_{\mathcal{M}}})$ reduces to a multivariate Gaussian, allowing for an analytical solution for the gradient.
+The specific discretization update rule for network parameters samples under functional stochastic gradient Langevin dynamics (fSGLD) is given by
+$\mathbf{w}_{t+1}=$ $\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f_{t}\Big{[}-\frac{N}{n}\sum_{i=1}^{n}\nabla_{f}\log p(y_{i}|f_{t}(x_{i};\mathbf{w}_{t}))-$ (6)
+$\hskip 62.59596pt\nabla_{f^{\mathbf{X}_{\mathcal{M}}}}\log P_{0}(f_{t}^{\mathbf{X}_{\mathcal{M}}})\Big{]}+\sqrt{2\epsilon_{t}}\eta_{t}.$
+Like the standard SGLD, with the discretization step-size $\epsilon_{t}$ decays to zero, the stationary measure of fSGLD continues to be the target posterior $P_{f|\mathcal{D}}$ over functions.
+Note that in contrast to the naive parameter-space SGLD in Equation 1, which is driven solely by limited parameter information, the discretization update rule for samples of network parameters in Equation 6 under our fSGLD scheme incorporates information about the function mapping.
+This additional functional information plays a crucial role in achieving more accurate posterior inference.
+The pseudocode for fSGLD is shown in Algorithm 1 in Appendix D.
+3.2 Functional Hamiltonian dynamics for BNNs
+For the naive Hamiltonian dynamics, $X=(\mathbf{w},\mathbf{z})$ represents the augmentation of model parameters.
+We now let $X=(f,g)$ and define the Hamiltonian dynamics in function space as $H(f,g)=U(f)-\log p(g)$, where $p(g)$ is the functional auxiliary probability measure induced by the distribution of auxiliary variable $\mathbf{z}$.
+We then design the functional Hamiltonian dynamics for $f(\cdot;\mathbf{w})$ and $g(\cdot;\mathbf{z})$ as
+$\mathrm{d}f_{t}(\cdot;\mathbf{w})=-(\nabla_{\mathbf{w}}f_{t})^{T}\nabla_{\mathbf{z}}g_{t}\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (7)
+$\mathrm{d}g_{t}(\cdot;\mathbf{z})=-(\nabla_{\mathbf{z}}g_{t})^{T}\nabla_{\mathbf{w}}f_{t}\nabla_{f}U(f_{t})\mathrm{d}t.$
 Table 1: Drift term and diffusion coefficient for naive SGMCMC and functional SGMCMC algorithms.
@@
 Proposition 3.2.
-The (marginal) stationary distribution of the functional Hamiltonian dynamics defined in Equation 10 is the target posterior distribution over functions $p(f|\mathcal{D})$.
+The (marginal) stationary probability measure of the functional Hamiltonian dynamics defined in Equation 7 is the target posterior over functions $P_{f|\mathcal{D}}$.
 Proof.
-The proof is given in Appendix B. ∎
-Similar to the fSGLD, we also transform the above function-space Hamiltonian dynamics to its corresponding weight-space dynamics of $\mathbf{w}$ and $\mathbf{z}$ as follows:
-$\mathrm{d}\mathbf{w}_{t}$ $=\frac{\partial H}{\partial\mathbf{z}}=-\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (11)
-$\mathrm{d}\mathbf{z}_{t}$ $=-\frac{\partial H}{\partial\mathbf{w}}=-\nabla_{\mathbf{w}}f\cdot\nabla_{f}U(f_{t})\mathrm{d}t,$
+The proof is given in Appendix C. ∎
+Similar to the fSGLD, we also transform the above function-space Hamiltonian dynamics to its corresponding parameter-space dynamics of $\mathbf{w}$ and $\mathbf{z}$ as follows:
+$\mathrm{d}\mathbf{w}_{t}$ $=\frac{\partial H}{\partial\mathbf{z}}=-\nabla_{\mathbf{z}}g_{t}\cdot\nabla_{g}\log p(g_{t})\mathrm{d}t,$ (8)
+$\mathrm{d}\mathbf{z}_{t}$ $=-\frac{\partial H}{\partial\mathbf{w}}=-\nabla_{\mathbf{w}}f_{t}\cdot\nabla_{f}U(f_{t})\mathrm{d}t,$
 where the random diffusion term is 0 for the full-batch version.
 For the stochastic gradient version, the discretization update rule for samples of $\mathbf{w}$ and $\mathbf{z}$ under the functional stochastic gradient Hamiltonian dynamics (fSGHMC) is as follows:
-$\mathbf{w}_{t+1}=$ $~{}\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})$ (12)
-$\mathbf{z}_{t+1}=$ $~{}\mathbf{z}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f\cdot\nabla_{f}\tilde{U}(f_{t})$
-$\hskip 14.22636pt+\epsilon_{t}C\nabla_{\mathbf{z}}g\cdot\nabla_{g}\log p(g_{t})+\sqrt{2C\epsilon_{t}}\eta_{t}.$
-The stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$ (see Appendix B).
-Like the discussion in the last section, the diffusion process for $\mathbf{w}$ and $\mathbf{z}$ under fSGHMC scheme in Equation 12 containing abundant functional information is obviously different from the naive parameter-space SGHMC in Equation 2.
-The pseudocode for fSGHMC is shown in Algorithm 2 in Appendix C.
-For our functional SGMCMC, the predictive distribution is obtained from the following integration process:
-$p(y^{*}|x^{*},\mathcal{D})$ $=\int p(y^{*}|f(x^{*}))p(f|\mathcal{D})\mathrm{d}f$ (13)
+$\mathbf{w}_{t+1}=$ $~{}\mathbf{w}_{t}-\epsilon_{t}\nabla_{\mathbf{z}}g_{t}\cdot\nabla_{g}\log p(g_{t})$ (9)
+$\mathbf{z}_{t+1}=$ $~{}\mathbf{z}_{t}-\epsilon_{t}\nabla_{\mathbf{w}}f_{t}\cdot\nabla_{f}\tilde{U}(f_{t})$
+$\hskip 14.22636pt+\epsilon_{t}C\nabla_{\mathbf{z}}g_{t}\cdot\nabla_{g}\log p(g_{t})+\sqrt{2C\epsilon_{t}}\eta_{t}.$
+The stationary probability measure of such functional stochastic gradient Hamiltonian can still be guaranteed as the target functional posterior $P_{f|\mathcal{D}}$ (see Appendix C).
+Like the discussion in the last section, the diffusion process for $\mathbf{w}$ and $\mathbf{z}$ under fSGHMC scheme in Equation 9 incorporates abundant functional information is obviously different from the naive parameter-space SGHMC in Equation 2.
+The pseudocode for fSGHMC is shown in Algorithm 2 in Appendix D.
+For our functional MCMC schemes, the predictive distribution is obtained from the following integration process:
+$p(y^{*}|x^{*},\mathcal{D})$ $=\int p(y^{*}|f(x^{*}))P_{f|\mathcal{D}}\mathrm{d}f$ (10)
 $\approx\frac{1}{S}\sum_{j=1}^{S}p(y^{*}|f(x^{*};\mathbf{w}^{(j)})),$
-where function draws of posterior over functions are generated from corresponding samples of weight $\mathbf{w}^{(j)}(j=1,2,...,S)$ using fSGLD or fSGHMC.
+where function draws of posterior over functions are generated from corresponding samples of network parameters $\mathbf{w}^{(j)}(j=1,2,...,S)$ using fSGLD or fSGHMC.
 The comparisons of the definitions for the drift term and the diffusion coefficient of popular exiting SGMCMC algorithms and our functional fSGLD and fSGHMC are summarized in Table 1.
@@
 Functional Variational Inference for BNNs.
-With unresolved concerns about parameter-space weight prior to posterior approximation in BNNs, Bayesian inference in function space has attracted increasing attention.
-In order to specify meaningful functional prior like Gaussian processes, Sun et al.
-(2019) proposed a kind of functional variational Bayesian neural networks by maximizing the functional evidence lower bound (ELBO) to optimize the Kullback–Leibler (KL) divergence between the posterior processes and variational posterior over functions explicitly.
-They solved the KL divergence between infinite-dimensional stochastic processes on a finite marginal measurement set and used the Spectral Stein gradient estimator for implicit functional prior and variational posterior over functions of BNNs (Shi et al., 2018).
-However, Burt et al.
-(2020) pointed out some limitations of employing KL divergence in infinite-dimensional function space.
-The variational posterior should be absolutely continuous with respect to the prior measure to guarantee the existence of the Radon-Nikodym derivatives involved in the definition of functional KL divergence (Matthews et al., 2016; Wild et al., 2022).
-Otherwise, for example, the KL divergence between distributions over functions generated from two BNNs with different network architectures is ill-defined and could be infinite (Ma & Hernández-Lobato, 2021).
-To avoid this issue, Rudner et al.
-(2022) proposed to approximate the distributions over functions of a BNN prior and variational posterior that share the same network structures as two Gaussian distributions via the linearization of their mean parameters and derived a tractable and well-defined functional variational objective for BNNs.
-Considering the potential weaknesses of KL divergence for stochastic processes, Tran et al.
-(2022) proposed to match a BNN prior to a GP prior by minimizing the 1-Wasserstein distance (Kantorovich, 1960) to obtain a more interpretable functional prior for BNNs.
+To specify meaningful functional priors like Gaussian processes, Sun et al., (2019) proposed a functional evidence lower bound (ELBO) to optimize the Kullback–Leibler (KL) divergence between variational posterior and the true posterior in function space explicitly.
+They solved the KL divergence between infinite-dimensional stochastic processes on finite marginal measurement sets.
+Due to the analytical intractability of the supremum over marginal KL divergences in functional ELBO, Rudner et al., (2022) proposed to approximate the posterior and prior over functions of BNNs as Gaussian distributions through local linearization and derived a more tractable functional variational objective for BNNs.
+Considering the potential weaknesses of KL divergence for stochastic processes Burt et al., (2020), Tran et al., (2022) proposed to match a BNN prior to a GP prior by minimizing the 1-Wasserstein distance (Kantorovich,, 1960) to obtain a more interpretable functional prior.
 Markov Chain Monte Carlo for BNNs.
-Papamarkou et al.
-(2022) summarized the challenges in applying MCMC approaches to posterior inference for BNNs in four aspects: computational cost, model structure, weight symmetries, and prior specification.
-MCMC methods have not been broadly adopted in posterior inference for large modern BNNs until the introduction of scalable Monte Carlo algorithms, such as the stochastic gradient MCMC (Welling & Teh, 2011; Chen et al., 2014; Ahn et al., 2014; Garriga-Alonso & Fortuin, 2021).
-Ma et al.
-(2015) designed a unifying framework for stochastic gradient MCMC whereby all past dynamics-based samplers and their minibatch variants can be cast into it.
-For the multimodal properties of posterior, Zhang et al.
-(2020) proposed a cyclical step size schedule in stochastic gradient MCMC for BNNs to improve the sampling efficiency by which they proved the non-asymptotic convergence of their method.
-On the other hand, Cobb & Jalaian (2021) emphasized the limitations of the use of stochastic gradients in HMC (Shahbaba et al., 2014), and proposed a scaling full HMC for BNNs by a symmetric splitting integration scheme.
-And Izmailov et al.
-(2021) also used full-batch HMC to investigate the foundational properties of posterior for BNNs.
-Alexos et al.
-(2022) proposed a structured stochastic gradient MCMC that combines structure variational inference and MCMC to improve the slow mixing in high dimensional BNNs.
-(a) fSGLD
-(b) fSGHMC
-(c) FBNN
-(d) SGLD
-(e) SGHMC
-(f) IFBNN
-Figure 1: 1-D extrapolation example.
-The green line is the ground true function, and the blue lines correspond to the mean of samples from posterior predictions.
-Black dots denote 20 training points; shadow areas represent the predictive standard deviations.
+Papamarkou et al., (2022) summarized the challenges in applying MCMC approaches to posterior inference for BNNs in four aspects: computational cost, model structure, weight symmetries, and prior specification.
+MCMC methods had limited success in being broadly adopted for posterior inference in large modern BNNs until the introduction of scalable Monte Carlo algorithms, such as the stochastic gradient MCMC (Welling and Teh,, 2011; Chen et al.,, 2014; Ahn et al.,, 2014; Garriga-Alonso and Fortuin,, 2021).
+Ma et al., (2015) designed a unifying framework that casts all existing dynamics-based samplers and their minibatch variants into it.
+Regarding the multimodal properties of posterior distributions, Zhang et al., (2020) proposed a cyclical step-size schedule in stochastic gradient MCMC for BNNs to improve the sampling efficiency.
+On the other hand, Cobb and Jalaian, (2021) emphasized the limitations of stochastic gradients in HMC and proposed a scaling full-batch HMC for BNNs using a symmetric splitting integration scheme.
+However, all these MCMC methods are performed in parameter space.
 5 Experiments
-We evaluate the predictive performance and uncertainty quantification of our functional stochastic gradient MCMC methods on several benchmark tasks, including a synthetic extrapolation example, multivariate regressions on UCI datasets, image classification tasks, and contextual bandits.
-We compare fSGLD and fSGHMC with naive SGLD, SGHMC, and competing functional variational inference methods to illustrate the sampling efficiency of our functional stochastic gradient MCMC methods.
+We evaluate the predictive performance and uncertainty quantification of our functional stochastic gradient MCMC schemes on several benchmark tasks, including a synthetic extrapolation example, multivariate regressions on UCI datasets, image classification tasks, and contextual bandits.
+We compare fSGLD and fSGHMC with naive SGLD, SGHMC, and competing functional variational inference methods to illustrate the sampling efficiency of our functional stochastic gradient MCMC schemes.
 5.1 Extrapolation on Synthetic Data
-To evaluate the fitting ability and uncertainty estimation of our method, we first consider a 1-D oscillation curve extrapolation on a synthetic dataset.
+To evaluate the fitting ability and uncertainty estimation, we first consider a 1-D oscillation curve extrapolation on a synthetic dataset.
 For input, we randomly sampled half of 20 observation points from Uniform$(-0.75,-0.25)$, and the other half are from Uniform$(0.25,0.75)$.
 The output is modeled through the polynomial function: $y=\sin(3\pi x)+0.3\cos(9\pi x)+0.5\sin(7\pi x)+\epsilon$ with noise $\epsilon\sim\mathcal{N}(0,0.5^{2})$.
-We compared our fSGLD and f SGHMC with naive SGLD, SGHMC, and two representative functional variational methods: FBNN (Sun et al., 2019) and IFBNN (Wu et al., 2023).
+We compared our fSGLD and fSGHMC with naive SGLD, SGHMC, and two representative functional variational methods: FBNN (Sun et al.,, 2019) and IFBNN (Wu et al.,, 2023).
 The model is a fully-connected neural network with two hidden layers.
@@
 We found that FBNN in Figure 1(c) is unable to fit the target curve in the training data range $[-0.75,0.25]$ and performs poorly for uncertainty quantification.
-Figure 1(f) shows that IFBNN also severely underestimates the uncertainty, which is consistent with the findings of Ormerod & Wand (2010) and Zhang et al.
-(2018) that variational approximation typically underestimates the posterior variance.
+Figure 1(f) shows that IFBNN also severely underestimates the predictive uncertainty, which is consistent with the findings of Ormerod and Wand, (2010) and Zhang et al., (2018) that variational approximation typically underestimates the posterior variance.
+More analysis of mixing rate and computational complexity are shown in Appendix F and Appendix G, respectively.
+Appendix H presents a wide ablation study about the sample size, measurement points and functional prior.
+(a) fSGLD
+(b) fSGHMC
+(c) FBNN
+(d) SGLD
+(e) SGHMC
+(f) IFBNN
+Figure 1: 1-D extrapolation example.
+The green line is the ground true function, and the blue lines correspond to the mean of samples from posterior predictions.
+Black dots denote 20 training points; shadow areas represent the predictive standard deviations.
+For more details, see Appendix E.
 5.2 UCI Regression
-In this experiment, we evaluate our method for multivariate regression tasks on 7 benchmark UCI datasets: Yacht, Boston, Concrete, Energy, Wine, Kin8nm and Protein.
-We use a 2- hidden-layer fully connected neural network.
+In this experiment, we evaluate our methods on multivariate regression tasks across seven benchmark UCI datasets: Yacht, Boston, Concrete, Energy, Wine, Kin8nm and Protein.
+We use a two-hidden-layer fully connected neural network.
 For all sampling methods, we run 500 iterations for the burn-in stage and collect 15 samples.
 Performance is measured via average root mean square error (RMSE) and the test negative log-likelihood (NLL).
-For each dataset, we construct 10 random 90-to-10 train-test splits and report the mean and standard deviation of RMSE and NLL over the splits in Table 2 and Table 3, respectively.
+For each dataset, we construct 10 random 90-to-10 train-test splits and report the mean and standard deviation of RMSE and NLL over these splits in Table 2 and Table 3, respectively.
 fSGHMC outperforms all baselines for RMSE on all datasets, and fSGLD also achieves better accuracy values than naive SGMCMC methods and functional variational inference, demonstrating superior performance of our functional SGMCMC in predictive accuracy.
-Moreover, fSGHMC and fSGLD obtain better NLL results on most of the datasets (5/7), showing that our method is also very competitive in uncertainty quantification.
+Moreover, fSGHMC and fSGLD obtain better NLL results on most of the datasets (5/7), showing that our methods are also very competitive in uncertainty quantification.
 Table 2: The table shows the results of average RMSE for multivariate regression on UCI datasets.
 We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
+Bold indicates statistically significant best results ($p<0.01$ with t-test).
 RMSE
@@
 We split each dataset randomly into 90% training data and 10% test data, and this process is repeated 10 times to ensure validity.
+Bold indicates statistically significant best results ($p<0.01$ with t-test).
 NLL
@@
 5.3 Imagine Classification and OOD Detection
-In this section, we investigate the scalability of functional SGMCMC on image classification tasks with high-dimensional inputs.
-We evaluate the in-distribution predictive performance and out-of-distribution (OOD) detection ability on MNIST (LeCun et al., 2010), FashionMNIST(Xiao et al., 2017) and CIFAR-10(Krizhevsky et al., 2009).
-We compare our fSGLD and fSGHMC with naive SGLD, SGHMC, functional variational FBNN and IFBNN.
-We use ResNet-18 architecture (He et al., 2016) for all methods on these three datasets, and the batch size is 125.
+In this section, we investigate the scalability of our methods on image classification tasks with high-dimensional inputs.
+We evaluate the in-distribution predictive performance and out-of-distribution (OOD) detection ability on MNIST (LeCun et al.,, 2010), FashionMNIST(Xiao et al.,, 2017) and CIFAR-10(Krizhevsky et al.,, 2009).
+We use ResNet-18 architecture (He et al.,, 2016) for all methods on these three datasets, and the batch size is 125.
 For all sampling methods, we run for 60 burn-in iterations and collect 10 samples on the MNIST dataset, 100 burn-in iterations and collect 10 samples on FashionMNIST, 80 burn-in iterations and draw 10 samples on CIFAR10.
-The functional GP prior with the RBF kernel is pre-trained on the test dataset for all functional methods.
 We randomly sample 40 points from training data in every iteration as the measurement points for the gradient estimation of log functional prior distribution in fSGLD and fSGHMC.
@@
 In such problems, an agent interacts with an unknown environment repeatedly and chooses to take an optimal action in each round of interaction.
-Thompson sampling (Thompson, 1933; Russo & Van Roy, 2016) is a widely used algorithm for exploration strategies in contextual bandits, it (i) first samples a current posterior to obtain a model configuration; (ii) then adaptively chooses an optimal action based on the current context under the sampled model configuration and observes the corresponding reward; (iii) updates the posterior based on the context, action and reward tuples in this interaction round.
+Thompson sampling (Thompson,, 1933; Russo and Van Roy,, 2016) is a widely used algorithm for exploration strategies in contextual bandits, it (i) first samples a current posterior to obtain a model configuration; (ii) then adaptively chooses an optimal action based on the current context under the sampled model configuration and observes the corresponding reward; (iii) updates the posterior based on the context, action and reward tuples in this interaction round.
 This section evaluates the ability to guide exploration on the UCI Mushroom dataset with 8124 instances.
@@
 An agent can observe mushroom features as the context in each interaction and choose to eat or reject a mushroom.
-We compare our fSGLD with several baselines, including naive SGLD, functional variational FBNN and IFBNN, parametric-space variational method Bayes By Backprop (BBB) (Blundell et al., 2015).
-We follow the basic settings by Blundell et al.
-(2015) and consider three different reward patterns: if the agent eats an edible mushroom, it receives a reward of 5 and a reward of 0 if the agent chooses to reject.
+We compare our fSGLD with several baselines, including naive SGLD, functional variational FBNN and IFBNN, parametric-space variational method Bayes By Backprop (BBB) (Blundell et al.,, 2015).
+We follow the basic settings by Blundell et al., (2015) and consider three different reward patterns: if the agent eats an edible mushroom, it receives a reward of 5 and a reward of 0 if the agent chooses to reject.
 On the other hand, if the agent eats a poisonous mushroom, it receives a reward of -35 with probabilities of 0.4, 0.5, and 0.6, respectively, for three different patterns.
@@
 6 Discussion and Conclusions
-In this paper, we have lifted the Itô diffusion of network weights in BNNs to function space and proposed a new functional SGMCMC scheme for BNNs.
-We have designed the functional Langevin dynamics and Hamiltonian dynamics for the function mapping defined by BNNs directly, which can incorporate informative functional prior into posterior inference.
-Moreover, we proved that the stationary distribution of our functional SGMCMC is the posterior distribution over functions.
-Empirically, we demonstrated that our functional SGMCMC approaches can leverage functional priors to yield high predictive performance and principled uncertainty quantification on several benchmark tasks.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Further Background
-Itô Lemma (Itô, 1951) is a fundamental result in stochastic calculus to find the differential of a function of a stochastic process, which serves as the stochastic calculus counterpart of the chain rule.
+In this paper, we lift the Itô diffusion of network parameters in BNNs onto function space and propose novel functional MCMC schemes for BNNs.
+Specifically, we design functional Langevin dynamics and functional Hamiltonian dynamics for the function mapping defined by BNNs and derive corresponding tractable sampling methods in parameter space, which can incorporate informative functional prior into posterior inference.
+Moreover, we proved that the stationary measure of our functional dynamics is the posterior over functions.
+Empirically, we demonstrate that our functional MCMC schemes effectively leverage functional priors to yield superior predictive performance and principled uncertainty quantification on several benchmark tasks.
+Appendix A Notation Table
+Table 5 is the notation table to demonstrate the notation used in this paper.
+Table 5: Notation table
+Notation Meanings
+$\mathcal{D}=\left\{\mathbf{X}_{\mathcal{D}},\mathbf{Y}_{\mathcal{D}}\right\}$ Training dataset
+$\mathcal{X}\subseteq\mathbb{R}^{p}$ ($p$-dimensional) input space
+$\mathcal{Y}\subseteq\mathbb{R}^{c}$ ($c$-dimensional) output space
+$(\Omega,\mathcal{A},P)$ Probability space on $\mathbb{R}^{k}$
+$\mathbb{H}$ Infinite-dimensional function space (Polish space)
+$\mathcal{B}(\mathbb{H})$ Borel $\sigma$-algebra on $\mathbb{H}$
+$\mathcal{P}(\mathbb{H})$ The space of Borel probability measures on $\mathcal{B}(\mathbb{H})$.
+$X$ Stochastic process
+$\mathbf{X}_{\mathcal{M}}$ Finite measurement points from input space
+$\mathbf{w}\in\mathbb{R}^{k}$ Random network parameters
+$\mathbf{z}\in\mathbb{R}^{k}$ Momentum variable
+$\epsilon$ Discretization step size
+$\eta$ Standard Gaussian noise
+$f(\cdot;\mathbf{w})$ Random function mapping defined by a BNN parameterized by $\mathbf{w}$
+$p_{0}(\mathbf{w})$ Prior distribution over network parameters
+$p(\mathbf{w}|\mathcal{D})$ Posterior distribution over network parameters
+$p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$ Likelihood function evaluated on the training data
+$p(\mathbf{z})$ Auxiliary probability distribution
+$P_{0}$ Functional Prior measure
+$\operatorname{supp}(P_{0})$ Topological support of $P_{0}$
+$E$ A non-empty subset of $\operatorname{supp}(P_{0})$
+$P_{f|\mathcal{D}}$ Posterior measure over functions
+$\Phi(f)$ Negative log-likelihood as $-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$
+$p(g)$ Functional auxiliary probability measure
+$U(\mathbf{w})$ Potential energy function in parameter-space Langevin dynamics
+$H(\mathbf{w},\mathbf{z})$ Potential energy function in parameter-space Hamiltonian dynamics
+$I_{0}(f)$ Onsager–Machlup (OM) functional for $P_{0}$
+$U(f)$ Potential energy functional in functional Langevin dynamics
+$H(f,g)$ Potential energy functional in functional Hamiltonian dynamics
+$\mu(\cdot)$ Drift term in the SDE of Itô diffusion
+$\sigma(\cdot)$ Diffusion coefficient in the SDE of Itô diffusion
+$B$ Standard Wiener process (Brownian motion)
+$\pi(\cdot)$ Stationary distribution (measure)
+Appendix B Further Background
+Itô diffusion
+Dynamics-based MCMC methods are rooted in the general framework of Itô diffusion (Øksendal,, 2003), which is commonly used to model the evolution of particles in a system.
+Given a stochastic process $X:[0,\infty)\times\Omega\rightarrow\mathbb{R}^{n}$ defined on a probability space $(\Omega,\Sigma,P)$, an Itô diffusion in n-dimensional Euclidean space, driven by the standard Wiener process (Brownian motion), satisfies the following specific type of stochastic differential equation (SDE):
+$\mathrm{d}X_{t}=\mu(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},$ (11)
+where $X_{t}$ represents the state of the stochastic process at time $t$, $\mu(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a vector field describing the deterministic drift term, and $\sigma(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$ is a matrix field denotes the diffusion coefficient for $X_{t}$.
+Both $\mu(\cdot)$ and $\sigma(\cdot)$ are assumed to satisfy the standard Lipschitz continuity condition (Ghosh,, 2010).
+The term $B\in\mathbb{R}^{n}$ refers to an n-dimensional Wiener process (Brownian motion), with $\mathrm{d}B_{t}$ being the increment of a Wiener process, distributed as $\mathrm{d}B_{t}\sim\mathcal{N}(0,\mathrm{d}t\cdot\mathbf{I}_{n})$.
+The probability density function $p(x,t)$ of $X_{t}$ is governed by the Fokker-Planck (FP) equation (Risken,, 1996):
+$\frac{\partial p(x,t)}{\partial t}=-\sum_{i}\frac{\partial}{\partial x_{i}}\left[\mu_{i}(X_{t})p(x,t)\right]+\sum_{i,j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\left[D_{i,j}(x,t)p(x,t)\right],$ (12)
+where $D_{i,j}(x,t)=\frac{1}{2}\sigma(X_{t})\sigma(X_{t})^{T}$.
+This equation serves as the foundation for deriving dynamics-based MCMC sampling methods.
+The diffusion process reaches a stationary state when the FP equation equals zero, then simulating this diffusion process is equivalent to sampling from the stationary distribution, denoted by $\pi(x)$.
+By carefully designing $\mu(\cdot)$ and $\sigma(\cdot)$, one can sample from the target distribution as the stationary distribution.
+For example, for the commonly used Langevin dynamics, where $\mu(X)=-\nabla U(X)$, $\sigma(X)=\sqrt{2}\cdot\mathbf{I}_{n}$, and $U(\cdot)$ is a potential energy function.
+The stationary distribution for $p(x,t)$ is given by the Boltzmann distribution $\pi(x)\propto\exp(-U(X))$.
+Thus, when sampling the posterior $p(\mathbf{w}|\mathcal{D})$ in BNNs, the potential energy function is given by $U(\mathbf{w})=-\log p(\mathbf{Y_{\mathcal{D}}}|\mathbf{X_{\mathcal{D}}};\mathbf{w})-\log p(\mathbf{w})$, and $\sigma(\mathbf{w})=\sqrt{2}\cdot\mathbf{I}_{k}$, ensuring that the target posterior as the stationary distribution.
+Itô Lemma
+Itô Lemma (Itô,, 1951) is a fundamental result in stochastic calculus to find the differential of a function of a stochastic process, which serves as the stochastic calculus counterpart of the chain rule.
 For an Itô diffusion, let $f(X_{t})$ be an arbitrary twice differentiable scalar function of real variables $X_{t}$, then the differential of $f(X_{t})$ can be derived from the Taylor series expansion of the function as
 $\mathrm{d}f(X_{t})=$ $\left(\mu(X_{t})\frac{\partial f}{\partial x}+\frac{\sigma^{2}(X_{t})}{2}\frac{\partial^{2}f}{\partial x^{2}}\right)\mathrm{d}t+\sigma(X_{t})\frac{\partial f}{\partial x}\mathrm{d}B_{t},$ (13)
-which implies that $f(X_{t})$ is itself an Itô diffusion (Brzeźniak et al., 2008).
-Appendix B Theoretical Proof
-The Fokker–Planck equation (Risken, 1996) for the corresponding evolution of the probability density $p(x,t)$ of a diffusion process $X_{t}$ is as
-$\frac{\partial p(x,t)}{\partial t}=-\sum_{i}\frac{\partial}{\partial x_{i}}\left[\mu_{i}(X_{t})p(x,t)\right]+\sum_{i,j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\left[D_{i,j}(x,t)p(x,t)\right],$ (15)
-where $D_{i,j}(x,t)=\frac{1}{2}\sigma(X_{t})\sigma(X_{t})^{T}$.
-When the Fokker–Planck equation is equal to zero, it means that the diffusion process has reached a stationary state and $p(x,t)$ is equal to the stationary distribution $\pi(X)$.
-Then, simulating this diffusion process is equivalent to sampling from the stationary distribution.
-The proof of Proposition 3.1 and Proposition 3.2 will be concise based on the general framework for dynamics-based sampling methods proposed by Ma et al.
-(2015).
-They proved that if $\mu(X_{t})$ and $\sigma(X_{t})$ are restricted to the following form
+which implies that $f(X_{t})$ is itself an Itô diffusion (Brzeźniak et al.,, 2008).
+Onsager–Machlup functional
+(see e.g., Lambley, (2023), Definition 2.4.) Suppose $\mathbb{H}$ is a separable metric space and $P_{0}\in\mathcal{P}(\mathbb{H})$.
+Assume that $\varnothing\neq E\subseteq\operatorname{supp}(P_{0})$, then a functional $I_{0}:E\rightarrow\mathbb{R}$ is an Onsager–Machlup (OM) functional for $P_{0}$ if for all $f,f^{\prime}\in E$,
+$\lim_{r\rightarrow 0}\frac{P_{0}(B_{r}(f)}{P_{0}(B_{r}(f^{\prime})}=\exp(I_{0}(f^{\prime})-I_{0}(f)),$ (14)
+where $B_{r}(\cdot)$ denotes the closed ball of radius $r$ in $\mathbb{H}$ and $\operatorname{supp}(P_{0}):=\{f\in\mathbb{H}\ |\ P_{0}(B_{r}(f))>0\ \text{for all}\ r>0\}$.
+Lambley, (2023) (in Proposition 2.6.) proved that posterior measure $P_{f|\mathcal{D}}\in\mathcal{P}(\mathbb{H})$ defined as $P_{f|\mathcal{D}}(\mathrm{d}f)\propto\exp(-\Phi(f))P_{0}(\mathrm{d}f)$ has OM functional $I^{y}:E\rightarrow\mathbb{R}$ given by $I^{y}(f)=I_{0}(f)+\Phi(f)$, where $\Phi(f)$ is the continuous potential, in essence the negative log-likelihood as $-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))$.
+OM functional can be interpreted heuristically as the negative logarithm of the Lebesgue density, but this interpretation cannot be taken literally in infinite-dimensional spaces, as there is no Lebesgue measure in such settings.
+For instance, the OM functional for a Gaussian measure on an infinite-dimensional Banach space is typically defined only on a specific subspace, known as the Cameron-Martin space, e.g., for a centred Gaussian measure, the OM functional is defined as $I:E\rightarrow\mathbb{R},I(f)=\frac{1}{2}||f||^{2}_{E}$.
+As a result, the OM functional does not need to be defined over the entire space $\mathbb{H}$.
+Appendix C Theoretical Proof
+The proof of Proposition 3.1 and Proposition 3.2 will be concise based on the general framework for dynamics-based sampling methods proposed by Ma et al., (2015).
+They proved in Theorem 1 that if $\mu(X_{t})$ and $\sigma(X_{t})$ are restricted to the following form
 $\mu(X)=$ $-[\mathbf{D}(X)+\mathbf{Q}(X)]\nabla G(X)+\Gamma(X),$ (15)
@@
 $\sigma(X)=$ $\sqrt{2\mathbf{D}(X)},$
-where $G(\cdot)$ is the potential energy function, $\mathbf{D}(\cdot)$ is a positive semidefinite matrix, and $\mathbf{Q}(\cdot)$ is a skew-symmetric curl matrix, then the Fokker–Planck equation can be transformed into a more compact form (Yin & Ao, 2006; Shi et al., 2012):
+where $G(\cdot)$ is the potential energy functional, $\mathbf{D}(\cdot)$ is a positive semidefinite matrix, and $\mathbf{Q}(\cdot)$ is a skew-symmetric curl matrix, then the Fokker–Planck equation can be transformed into a more compact form (Yin and Ao,, 2006; Shi et al.,, 2012) given by
 $\frac{\partial p(x,t)}{\partial t}=\nabla^{T}\cdot([\mathbf{D}(X)+\mathbf{Q}(X)][p(x,t)\nabla G(X)+\nabla p(x,t)]).$ (16)
-Then, it is easy to verify that the stationary distribution of the diffusion process is exactly as $\pi(X)\propto\exp(-G(X))$.
+Then, it is straightforward to verify that the stationary distribution of the diffusion process is exactly as $\pi(X)\propto\exp(-G(X))$.
 For the proof of Proposition 3.1, it is clear that the functional Langevin dynamics in Equation 3 can be cast into the above general framework as
-$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ $=[-(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)\nabla_{f}U(f_{t})+H_{\mathbf{w}}f]\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f)^{T}\mathrm{d}B_{t}$ (18)
+$\mathrm{d}f_{t}(\cdot;\mathbf{w})$ $=[-(\nabla_{\mathbf{w}}f_{t})^{T}(\nabla_{\mathbf{w}}f_{t})\nabla_{f}U(f_{t})+H_{\mathbf{w}}f]\mathrm{d}t+\sqrt{2}(\nabla_{\mathbf{w}}f_{t})^{T}\mathrm{d}B_{t}$ (17)
 $=\left(-[\mathbf{D}(f)+\mathbf{Q}(f)]\nabla_{f}U(f_{t})+\Gamma(f)\right)\mathrm{d}t+\sqrt{2\mathbf{D}(f)}\mathrm{d}B_{t},$
-where $\nabla_{f}U(f)=-\nabla_{f}\log p(\mathcal{D}|f)-\nabla_{f}\log p(f)$, $\mathbf{D}(f)=(\nabla_{\mathbf{w}}f)^{2}$, $\mathbf{Q}(f)=\mathbf{0}$, $\Gamma(f)=H_{\mathbf{w}}f$.
-Therefore, the Fokker-Planck equation of the probability density $p(f,t)$ for $f_{t}(\cdot;\mathbf{w})$ then can be derived as
+where $U(f)=-\log p(\mathbf{Y}_{\mathcal{D}}|f(\mathbf{X}_{\mathcal{D}};\mathbf{w}))+I_{0}(f)$ is an OM functional for $P_{f|\mathcal{D}}$, $\mathbf{D}(f)=(\nabla_{\mathbf{w}}f)^{T}(\nabla_{\mathbf{w}}f)$, $\mathbf{Q}(f)=\mathbf{0}$, $\Gamma(f)=H_{\mathbf{w}}f$.
+Therefore, the Fokker-Planck equation of the probability measure $p(f,t)$ for $f_{t}(\cdot;\mathbf{w})$ then can be derived as
 $\frac{\partial p(f,t)}{\partial t}=\nabla^{T}\cdot\left([\mathbf{D}(f)+\mathbf{Q}(f)][p(f,t)\nabla_{f}U(f)+\nabla p(f,t)]\right),$ (18)
-from which the stationary distribution of $p(f,t)$ can be verified as $\exp(-U(f))=p(f|\mathcal{D})$, that is, the target posterior distribution over functions.
+from which the stationary probability measure of $p(f,t)$ can be verified as $\exp(-U(f))=P_{f|\mathcal{D}}$, that is, the target posterior measure over functions.
 For the proof of Proposition 3.2, the functional Hamiltonian dynamics defined in Equation 7 can be written as
-$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&0\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}U(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t$ (20)
+$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g_{t}\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f_{t}&0\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}U(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t$ (19)
 $=-\mathbf{Q}(f,g)\nabla H(f_{t},g_{t})\mathrm{d}t,$
 where $\mathbf{Q}(f,g)$ is s a skew-symmetric curl matrix, $\mathbf{D}(f,g)=\mathbf{0}$.
-Therefore, the stationary distributions of the functional Hamiltonian dynamics defined in Equation 10 is $\pi(f,g)\propto\exp(H(f,g))$, and simply (marginalize) discard the auxiliary variables can obtain the target true posterior $p(f|\mathcal{D})$.
+Therefore, the stationary measures of the functional Hamiltonian dynamics defined in Equation 7 is $\pi(f,g)\propto\exp(H(f,g))$, and simply (marginalize) discard the auxiliary functions can obtain the target true posterior $P_{f|\mathcal{D}}$.
 For the stochastic gradient version, the functional stochastic gradient Hamiltonian dynamics (fSGHMC) for the discretization update rule for samples of $\mathbf{w}$ and $\mathbf{z}$ defined in Equation 9 is defined as follows:
-$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}\tilde{U}(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t+\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g)^{T}\end{array}\right]\mathrm{d}t,$ (21)
+$\mathrm{d}\left[\begin{array}[]{l}f_{t}(\cdot;\mathbf{w})\\ g_{t}(\cdot;\mathbf{z})\end{array}\right]$ $=-\left[\begin{array}[]{cc}0&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g_{t}\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f_{t}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g_{t}\end{array}\right]\left[\begin{array}[]{c}\nabla_{f}\tilde{U}(f_{t})\\ -\nabla_{g}\log p(g_{t})\end{array}\right]\mathrm{d}t+\left[\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\sqrt{2C}(\nabla_{\mathbf{z}}g_{t})^{T}\end{array}\right]\mathrm{d}t,$ (20)
 where $\mathbf{D}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$ and $\mathbf{Q}(f,g)=\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ is the same as that in full-batch functional Hamiltonian dynamics.
-Then, the (marginal) stationary distribution of such functional stochastic gradient Hamiltonian is still the target functional posterior $p(f|\mathcal{D})$.
+Then, the (marginal) stationary probability measure of such functional stochastic gradient Hamiltonian is still the target functional posterior $P_{f|\mathcal{D}}$.
 Moreover, the comparisons of matrix $\mathbf{D}(\cdot)$ and $\mathbf{Q}(\cdot)$ for naive SGMCMC methods and our functional SGMCMC are shown in Table 6.
@@
 fSGHMC $(f,g)$ $\tilde{H}(f,g)$   $\left(\begin{array}[]{cc}\mathbf{0}&-(\nabla_{\mathbf{w}}f)^{T}\nabla_{\mathbf{z}}g\\ (\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{w}}f&\mathbf{0}\end{array}\right)$ $\left(\begin{array}[]{cc}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&C(\nabla_{\mathbf{z}}g)^{T}\nabla_{\mathbf{z}}g\end{array}\right)$
-Appendix C Pseudocode for Functional SGMCMC.
+Appendix D Pseudocode for Functional SGMCMC.
 Algorithm 1 presents the pseudocode for fSGLD. And the pseudocode for fSGHMC is shown in Algorithm 2.
@@
   end for
-Appendix D Ablation Study
-D.1 The effects of the sample size for naive SGMCMC and functional SGMCMC
+Appendix E Experimental Setting
+Extrapolation on Synthetic Data
+In this experiment, we use $2\times 100$ fully connected tanh neural networks for all models.
+The functional GP prior with the RBF kernel is pre-trained on the 20 training points for 100 epochs for all functional methods.
+we use a decreasing step-size schedule for $\epsilon_{t}$ with 1e-3 as the initial value for all sampling methods.
+The specific decreasing schedule is as every 5000 iterations, the decay factor is 0.9.
+Multivariate regression on UCI datasets
+In this experiment, we use two-hidden-layer fully connected neural networks and each layer with 10 hidden units for all models.
+The functional GP prior with the RBF kernel is pre-trained for 100 epochs.
+We run 500 iterations for the burn-in stage and collect 15 samples in the following 1500 iterations for all sampling methods.
+The initial $\epsilon_{t}$ is 1e-3, and the decay factor is 0.9.
+Functional variational inference methods are trained for 2000 epochs for fair comparison.
+Imagine Classification and OOD Detection
+In this experiment, we use ResNet-18 architecture for all methods.
+The functional prior is a Dirichlet-based GP designed for classification tasks and is pre-trained for 500 epochs.
+The initial $\epsilon_{t}$ is 1e-2, the decay period is 20, and the decay factor is 0.99.
+For all sampling methods, we run for 60 burn-in iterations and collect 10 samples in the following 20 iterations on the MNIST dataset, 100 burn-in iterations and collect 10 samples in the following 100 iterations on Fashion-MNIST, 80 burn-in iterations and draw 10 samples in the following 20 iterations on CIFAR10.
+Functional variational inference methods are trained for 80, 200, and 100 epochs on MNIST, FashionMNIST and CIFAR-10, respectively.
+Contextual Bandits
+In this experiment, we use fully connected neural networks with input-100-100-output architecture for all methods.
+The GP prior with RBF kernel is pre-trained on 1000 randomly sampled points from training data for 100 epochs.
+The initial $\epsilon_{t}$ is 1e-2, and the decay factor is 0.9.
+ALL models are trained using the last 4096 input-output tuples in the training buffer with a batch size of 64 and training frequency of 64 for each iteration.
+Appendix F Analysis of Mixing Time
+The trajectories of our fSGLD and the naive parameter-space SGLD for both the 1-D extrapolation and UCI regression (Yacht) experiments are shown in Figure 3.
+For the extrapolation example, we use a two-hidden-layer fully connected neural network with 100 hidden units in each layer (resulting in a 10401-dimensional parameter space).
+For the UCI regression task, we use a 2×10 fully connected network (141-dimensional parameter space).
+As illustrated in Figure 3, our fSGLD converges rapidly to the stationary measure—within 2000 iterations for the extrapolation example and 500 iterations for the UCI regression task.
+This convergence rate is comparable to that of the parameter-space SGLD, highlighting the rapid mixing rate of our method.
+(a) fSGLD-toy
+(b) fSGLD-uci
+(c) SGLD-toy
+(d) SGLD-uci
+Figure 3: Log-posterior probability versus the number of iterations.
+Appendix G Computational Complexity
+The runtime comparisons of a single iteration for all methods on the 1-D extrapolation, UCI regression (yacht) and image classification (MNIST) are provided in Table 7.
+For our fSGLD, fSGHMC, functional variational IFBNN and FBNN methods, the functional GP prior was pre-trained for 100, 100, and 500 epochs, respectively, in these three experiments, which added only an extra 1s for both the 1-D and UCI experiments, and 3s for the classification task.
+Compared to functional variational inference methods, our functional SGMCMC schemes exhibit significantly better computational performance, and are only slightly slower than the naive parameter-space SGMCMC.
+This demonstrates the computational efficiency of our method while maintaining the advantages of incorporating functional priors for improved inference.
+Table 7: Runtime comparison of single iteration for all methods (in seconds).
+Model 1-D extrapolation UCI Image classification
+SGLD 0.0016 0.0025 4
+fSGLD 0.0044 0.0045 5.83
+SGHMC 0.0023 0.02 10.95
+fSGHMC 0.0051 0.044 13.64
+IFBNN 0.0094 0.0075 6.81
+FBNN 0.129 0.114 23.25
+Appendix H Ablation Study
+H.1 The effects of the sample size for naive SGMCMC and functional SGMCMC
 In this section, we investigate the effect of sample size on the results of naive SGMCMC and our functional SGMCMC by varying the number of samples.
@@
 Boston -0.509 $\pm$ 0.129 -2.068 $\pm$ 0.159 -2.352 $\pm$ 0.100 -3.060 $\pm$ 0.284
-D.2 The effects of the number of measurement points on the gradient estimation of functional prior
+H.2 The effects of the number of measurement points on the gradient estimation of functional prior
 In the 1-D extrapolation example in Section 5.1, we sampled 40 measurement points from 20 training data and an additional 40 inducing points from Uniform $(-1,1)$ to estimate the gradient of log functional prior for fSGLD and fSGHMC.
@@
 Boston -2.277 $\pm$ 0.089 -2.607 $\pm$ 0.329
-D.3 The effects of the kernel function on the functional GP prior
-(a) fSGLD-RBF
-(b) fSGLD-Linear
-(c) fSGLD-Matern
-(d) fSGHMC-RBF
-(e) fSGHMC-Linear
-(f) fSGHMC-Matern
+H.3 The effects of the functional GP prior
+(a) GP prior-RBF
+(b) GP prior-Linear
+(c) GP prior-Matern
+(d) fSGLD-RBF
+(e) fSGLD-Linear
+(f) fSGLD-Matern
+(g) fSGHMC-RBF
+(h) fSGHMC-Linear
+(i) fSGHMC-Matern
 Figure 6: The effects of the kernel function on the functional GP prior. The text after the short line in each subheading represents the corresponding kernel functions.
 Functional Gaussian processes (GPs) priors are able to encode prior knowledge about function properties (e.g., periodicity and smoothness) through corresponding kernel functions.
-We used RBF kernel in the 1-D extrapolation experiment, which is suited for modeling polynomial functions.
-We now consider two other kernel functions of GP prior to our functional SGMCMC: Matern kernel and Linear kernel (unsuitable for modeling polynomial oscillatory curves).
+We used RBF kernel in the 1-D extrapolation experiment, which is suited for modelling polynomial functions.
+We now consider two other kernel functions of GP priors to our functional SGMCMC: Matern kernel and Linear kernel (unsuitable for modelling polynomial oscillatory curves).
 The results are shown in Figure 6.
-The results from the Linear kernel in Figure 5(b) and Figure 5(e) for fSGLD and fSGHMC respectively show that the fitting performance deteriorates significantly in the observation region and also demonstrate a strong linear trend (converging to a horizontal line) in the middle unseen region $[-0.25,0.25]$.
+The top row presents the GP prior predictions for RBF kernel, Linear kernel and Matern kernel, respectively.
+The results from the mismatched Linear kernel in Figure 6(e) and Figure 6(h) for fSGLD and fSGHMC, respectively, show that the fitting performance deteriorates in the observation region and also demonstrate a strong linear trend (converging to a horizontal line) in the middle unseen region $[-0.25,0.25]$ since the Linear GP prior prediction shows a completely linear trend and falls to fit the target function.
 The uncertainty intervals in the leftmost $[-1,-0.75]$ and rightmost $[0.75,1]$ non-observation regions deviate far from the true function.
-On the contrary, the results of the Matern kernel are hardly different from those of the RBF kernel, due to the fact that it also has a strong ability to describe polynomial functions.
+On the contrary, the results of the Matern kernel are hardly different from those of the RBF kernel due to the fact that it also has a strong ability to describe polynomial functions.
+This indicates that our method can effectively incorporate functional prior information into the posterior inference.
+We also plot the predictions of fSGLD and fSGHMC with no pre-trained GP priors in Figure 7.
+The leftmost columns, Figure 7(a) and Figure 7(c), show the results for fSGLD and fSGHMC with pre-trained GP priors, as described in the main paper.
+The rightmost columns, Figure 7(b) and Figure 7(d), display the predictions without pre-trained GP priors, using a fixed constant mean function $m(x)=0$ and a kernel function $k(x,x^{\prime})=1$.
+Despite not having pre-trained priors, the posterior results remain quite similar to those with pre-trained GP priors, indicating that our methods are robust to the absence of prior pre-training.
+In Table 12, we report the RMSE and NLL results of fSGLD and fSGHMC on the UCI regression (Yacht) dataset without pre-trained GP priors.
+Even though the fixed GP priors without pre-training show slight degradation, the performance remains competitive compared to other baseline methods.
+(a) fSGLD
+(b) fSGLD
+(c) fSGHMC
+(d) fSGHMC
+Figure 7: 1-D extrapolation example without prior pre-training. (a) and (c) are the results with pre-trained prior; (b) and (d) are results without pre-trained prior.
+Table 12: RMSE and NLL of fSGLD and fSGHMC on UCI regression (yacht) with no pre-trained GP priors.
+Model pre-trained GP fixed GP
+RMSE
+fSGLD 0.41 0.58
+fSGHMC 0.25 0.51
+NLL
+fSGLD -2.46 -1.65
+fSGHMC -3.37 -2.52
+Furthermore, we analyze the influence of different pre-training epochs for the GP priors by comparing results with 100, 20, and 10 pre-training epochs.
+Figure 8 provides a detailed comparison: the leftmost columns, Figure 8(a), Figure 8(d) and Figure 8(g) depict the predictions of the three different GP priors, while the middle and rightmost columns illustrate the fSGLD and fSGHMC posterior results from these priors, respectively.
+Interestingly, while the predictions from the three GP priors differ significantly, there is no noticeable difference in the corresponding fSGLD and fSGHMC posteriors.
+This observation highlights that the effectiveness of our methods does not strongly depend on the GP prior pre-training procedure, making them more flexible in practical applications.
+(a) GP-100 epochs
+(b) fSGLD
+(c) fSGHMC
+(d) GP-20 epochs
+(e) fSGLD
+(f) fSGHMC
+(g) GP-10 epochs
+(h) fSGLD
+(i) fSGHMC
+Figure 8: 1-D extrapolation example of different pre-training epochs for GP prior.
+The leftmost column shows the predictions of GP prior with three different pre-training epochs: 100, 20, and 10 epochs from top to bottom.
+The middle and the rightmost columns are the corresponding fSGLD and fSGHMC posteriors from these three GP prior, respectively.